### v1

# Select, then Balance: A Plug-and-Play Framework for Exogenous-Aware Spatio-Temporal Forecasting

Abstract.
Spatio-temporal forecasting aims to predict the future state of dynamic systems and plays an important role in multiple fields. However, existing solutions only focus on modeling using a limited number of observed target variables. In real-world scenarios, exogenous variables can be integrated into the model as additional input features and associated with the target signal to promote forecast accuracy. Although promising, this still encounters two challenges: the inconsistent effects of different exogenous variables to the target system, and the imbalance effects between historical variables and future variables. To address these challenges, this paper introduces ExoST, a novel framework for modeling exogenous variables in spatio-temporal forecasting, which follows a “select, then balance” paradigm. Specifically, we first construct a latent space gated expert module, where fused exogenous information is projected into a latent space to dynamically select and recompose salient signals via specialized sub-experts. Furthermore, we design a siamese network architecture in which recomposed representations of past and future exogenous variables are fed into dual-branch spatio-temporal backbones to capture dynamic patterns. The outputs are integrated through a context-aware weighting mechanism to achieve dynamic balance during the modeling process. Extensive experiments on real-world datasets demonstrate the effectiveness, generality, robustness, and efficiency of our proposed framework.
1. Introduction
Spatio-temporal forecasting aims to model the dynamic evolution of spatial entities over time by leveraging historical spatio-temporal sequences to predict future states over multiple time steps (Fang et al., 2025; Shao et al., 2023; Chen and Liang, 2025a, b). This task finds critical applications in domains such as traffic management (Xia et al., 2025; Avila and Mezić, 2020), environmental monitoring (Liang et al., 2023; Wang et al., 2020), and weather forecasting (Yang et al., 2024; Bi et al., 2023). In recent years, researchers have developed various spatio-temporal (ST) neural encoders (Jin et al., 2023b, a) by combining building blocks with different relational inductive biases, such as graph convolution operator (Ma et al., 2025; Yu et al., 2017; Li et al., 2017) for spatial dependencies and sequence operator (Wu et al., 2019; Guo et al., 2019) for temporal patterns. However, these approaches predominantly rely on modeling limited observations of target system variables, focusing mainly on encoding the spatial and temporal dynamics of target signals, while overlooking a widely available yet underutilized source of information in real-world forecasting systems: spatio-temporal exogenous variables.
Spatio-temporal exogenous variables refer to contextual information that is related to the prediction system but originates from external dynamic systems (Berkani et al., 2023). Such variables may include historical information that co-occur with the target system as well as prior information about future time points provided by external auxiliary forecasting systems. These factors serve as critical complementary signals in prediction tasks, helping to reduce uncertainty while offering interpretability (Huang et al., 2025; Das et al., 2023; Wang et al., 2024; Arango et al., 2025; Olivares et al., 2023; Giganti et al., 2024). As illustrated in Fig. 1 (a), consider the sensor network of air quality monitoring stations in New York City. For any given monitoring station, air quality forecasting typically focuses on modeling various indicators of pollutant emissions (e.g., NO2, CO2) to predict future air conditions. However, air quality at a particular location may also be significantly influenced by surrounding contextual information. For instance, historical traffic congestion data in the region (e.g., traffic flow) could lead to sudden spikes in pollutant levels, while stable weather conditions predicted for the future (e.g., wind speed) may indicate a risk of continued deterioration in air quality.
Despite the considerable potential of incorporating exogenous information, two major challenges persist in practice:
✦ Challenge I: Inconsistent Variable Effects: In real-world scenarios, exogenous variables often exert varying degrees of influence on the prediction target. As shown in Figure 1(b), different exogenous variables display distinct temporal dynamics and correlations with the target signal. Some variables (e.g., traffic flow) may have substantial predictive value, while others (e.g., traffic incidents) may be marginally relevant or even introduce noise. Fig. 2 (a) further illustrates this phenomenon by showing the Pearson correlation coefficients between various exogenous variables and the target variable across randomly selected stations in the AQI-19 dataset. Despite this variability, existing approaches often adopt a one-size-fits-all strategy (Arango et al., 2025; Giganti et al., 2024), directly concatenating exogenous variables with input features, without selectively modeling their individual contributions. This lack of mechanism-aware integration limits performance and can even degrade predictive accuracy.
✦ Challenge II: Imbalanced Type Effects: In addition to inter-variable disparities, there also exist imbalanced type differences between past and future exogenous variables from different temporal contexts. As shown in Fig. 1(c), exogenous information may stem from historically perceived signals (e.g., past traffic data) or from proactively forecasted future signals (e.g., weather predictions), leading to temporal distribution asymmetry and context-dependent impacts on prediction outcomes. Fig. 2 (b) clearly demonstrates the distinct distribution patterns of normalized values across different types of exogenous variables in the AQI-19 dataset. Existing methods typically adopt joint modeling strategies that fail to distinguish the distinct influences of these two types of exogenous variables on the target system (Wang et al., 2024; Das et al., 2023; Olivares et al., 2023; Zhou et al., 2025). Moreover, even under decoupled modeling settings, simple weighted fusion is insufficient to effectively balance their heterogeneous effects.
To address the aforementioned challenges, we propose a novel exogenous variable-enhanced spatio-temporal forecasting framework, ExoST, which is built upon a core “select, then balance” paradigam to systematically exploit and integrate the contributions of exogenous variables, thereby improving overall predictive performance. The framework adopts a modular and flexible design that allows seamless integration with a wide range of mainstream spatio-temporal backbone networks. Specifically, ❶ we introduce an adaptive selection mechanism through a latent-space gated expert module, which projects both endogenous and multi-source exogenous variables into a condition latent space and activates specialized expert sub-networks via a gating mechanism. This enables the model to automatically select and combine the most informative exogenous factors. The module exhibits strong representational capacity and adaptability, effectively capturing relevant signals while suppressing irrelevant or redundant information, thus alleviating inconsistencies caused by variable heterogeneity. To address the imbalance between types of exogenous variables, ❷ we design a dual-branch siamese network architecture that independently models past and future exogenous information. These recomposed representations are fed into dual versions of any given spatio-temporal backbone to learn evolutionary patterns, and a context-aware balancer is employed to dynamic fusion and produce the final prediction.
In summary, our key contributions are as follows:
Adaptive Selection Mechanism for Inconsistent Variable Effects: We propose a Latent Space Gated Expert Module that effectively handle the latent associations among multi-source exogenous variables. Through an expert selection mechanism, it identifies and emphasizes the most predictive external signals, enhancing the quality of the semantic representation input.
Balance Modeling Strategy for Imbalanced Type Effects: We design a Siamese network architecture that separately models past and future exogenous information branch, enabling dynamic prediction balancing via a context-aware fusion gating mechanism.
A Flexible Plug-and-Play Framework: ExoST provides a general and flexible enhancement framework for spatio-temporal forecasting. It can be broadly applied to various backbones and consistently outperforms existing methods across multiple benchmarks, demonstrating both practical value and scalability.
2. Preliminaries
2.1. Related Work
2.1.1. Spatio-Temporal Forecasting.
Predicting future spatio-temporal signals from historical observations, a spatially extension of multivariate time series forecasting, has evolved significantly. Early statistical (Box and Pierce, 1970; Biller and Nelson, 2003) or independent time series models (Nie et al., 2023; Wu et al., 2022; Qiu et al., 2025; Seo and Lim, 2025; Fan et al., 2025) at each location proved insufficient for complex spatio-temporal dynamics (Shao et al., 2024), which involve intricate spatial correlations and heterogeneity (Chen and Liang, 2025a, b). This led to hybrid architectures (Cao et al., 2025; Kong et al., 2024; Ma et al., 2025; Liu et al., 2023; Shao et al., 2022; Wu et al., 2019; Guo et al., 2019; Li et al., 2017) that integrate spatial (e.g., graph operators for node interactions) and temporal (e.g., sequence operators for long-term trends) modeling to jointly learn rich spatio-temporal representations for accurate prediction. However, a critical limitation persists: these models largely focus on the system’s internal dynamics while neglecting exogenous variables—external inputs that are independent of the system’s state but directly influence its evolution, and whose inclusion is vital for robust and accurate predictions.
2.1.2. Exogenous Variables Modeling.
In practical time series forecasting scenarios, exogenous information is already integrated into the model as additional input features and concatenated with the target signal, enabling the model to learn the correlation between external factors and the target variable. Earlier studies explored statistical methods such as ARIMAX (Williams, 2001) and SARIMAX (Vagropoulos et al., 2016) that can understand the relationship between exogenous and endogenous sequences and autoregression. Advanced deep learning models discuss how to use future values of exogenous variables (Das et al., 2023; Olivares et al., 2023) and how to adapt to time lags and missing exogenous variable data scenarios (Wang et al., 2024; Zhou et al., 2025). In addition, some works (Huang et al., 2025; Liu et al., 2024; Jin et al., [n. d.]) also discuss how to use language models to align the semantic information of exogenous variables for enhancement. Limited works have explored the integration of exogenous variables in spatio-temporal forecasting scenarios, but they only focus on specific task scenarios and model structure design (Giganti et al., 2024; Dong et al., 2023; Zagouras et al., 2015; Lindström et al., 2014). In contrast, we systematically explore how to effectively model exogenous variables in spatio-temporal forecasting scenarios and adaptively weigh the impact of different exogenous variables.
2.2. Problem Formulation
Definition 1.
(Spatio-Temporal Data). Spatio-temporal data refers to observational records that capture how dynamic systems evolve in space and time. Each data point is associated with a specific geographic location and a specific timestamp. This type of data can usually be represented as a temporal sequence , where each is a snapshot of the spatial observation at time . The spatial is a sensor network consisting of discrete nodes, and each node has features at time (e.g., pollutant indicators CO2, NO2, etc. from a dynamic air monitoring systems.).
Definition 2.
(Exogenous Variables). Exogenous variables are external system factors that are independent of the spatio-temporal forecasting and are free of the internal dynamics. During the modeling process, these variables are fed into the model as known inputs (whether actual observations or predicted values). Typically, these data hold the same structure as spatio-temporal data, denoted as , and can be further subdivided into past exogenous variables (e.g., traffic flow and traffic speed at past moments) and future exogenous variables (e.g., future temperature and precipitation forecasts) according to their time attributes.
Problem 0.
(ST Forecasting with Exogenous Variables). Our goal is to learn an optimal prediction model for predicting future spatio-temporal signals from historical observations, optionally using exogenous inputs (and an optional graph structure prior ). Let be a training dataset sampled from a spatio-temporal process, where represents the features of nodes for the past time steps, and represents the target for the future time steps. Traditional spatio-temporal prediction optimizes the model parameters by minimizing the expected prediction loss:
where is the prediction model and measures the prediction error. When exogenous input is available, the goal becomes:
where and denote the exogenous inputs in the past and future, respectively. and are the corresponding feature dimensions.
3. Methodology
3.1. Overview
The ExoST is built on the “select-then-balance” paradigm, as shown in Figure 3. The process consists of two consecutive stages, which are used to address the two core challenges in exogenous-augmented spatio-temporal forecasting mentioned above: (i) inconsistent variable effects and (ii) unbalanced type effects.
Select stage (§3.2). A lightweight Latent-Space Gated Expert module projects heterogeneous exogenous variables into a condition latent manifold, and then activates a Mixture-of-Experts Selector to dynamically re-weight and re-compose the most meaningful signals. The output is two refined, exogenous-aware representations and , one for each type.
Balance stage (§3.3). A Siamese dual-branch architecture processes the two representations in parallel through the existing spatio-temporal backbones, and a Context-Aware Balancer adaptively balances their contributions based on instantaneous context. The final forecast is produced by a read-out layer on top of the fused hidden state.
Both stages are fully differentiable and backbone-agnostic: any modern ST encoder can be plugged in without architectural change.
3.2. Select: Latent-Space Gated Expert Module
The selection stage aims to address inconsistent variable effects by extract compact, adaptive semantic representations from input endogenous variables and heterogeneous, multi-source exogenous variables, laying the foundation for the subsequent balancing stage. This requires a carefully designed two-step process: (i) semantic alignment of diverse exogenous and endogenous variables in a unified semantic space via conditional embeddings; (ii) adaptive selection through a mixture-of-experts selector that reweights and combines existing features to mitigate antagonism caused by inconsistent variable effects.
3.2.1. Conditional Embedding
Conditional embeddings specifically address the semantic space alignment of endogenous variables with two types of exogenous variables. For each exogenous type (i.e., past, future), we first project the raw exogenous tensor and the target history input into a condition latent space via:
where is implemented as affine-transformation block:
where , are learned projections that align heterogeneous feature dimensions into a latent semantic space. Specifically, consistent with common settings in existing literature (Shao et al., 2024; Jin et al., 2024), the lengths of past and future steps are typically uniform. Therefore, to maintain this setting, we perform a direct element-wise add fusion. For more arbitrary cases, zero-padding preprocessing trick can also easily ensure alignment. Subsequently, non-linear activation functions project endogenous variables into latent conditional semantic spaces, derived from different types of exogenous variables, yielding conditional representations and , respectively.
3.2.2. Mixture-of-Experts Selector
As discussed, inconsistent variable effects between exogenous and endogenous variables can induce antagonistic or synergistic semantic degeneration. This manifests as channel entanglement in the derived conditional representations, where semantic mixing may lead to rank-deficient representations, complicating the learning process. To address this, we attach a mixture-of-experts gating (Jacobs et al., 1991; Shazeer et al., 2017; Fedus et al., 2022) module on top of each . Concretely, we maintain expert projections and a gating network parameterized by a single linear layer followed by function. For each conditional representation input, the gating score is applied at different time steps and locations and is calculated as follows:
where is projection layer projection layer, along with a normalized activation function, calculates each expert’s weight. This controls the weighting of each expert for the current input, simulating various semantic combinations. The final representation is then obtained through weighted combination:
This strategy explicitly disentangles representations using a latent bottleneck and selective recombination, rather than forcing the model to learn entangled mappings. This latent space gating method allows us to model multiple interactions between exogenous variables and the target representation within the fused conditional space, thereby adaptively acquiring suitable representation input and for the backbone network.
3.3. Balance: Dual-Branch Adaptive Fusion
The balancing stage aims to address the unbalanced type effect, that is, how to properly process the two semantic representations in the previous stage. This requires appropriately learning spatio-temporal dynamic patterns and effectively fusing them to achieve the final prediction. However, a key challenge arises from the inherent difference in temporal distribution between past and future exogenous information, complicating their effective fusion. To address this, we designed two crucial steps: (i) independent extraction of dynamic spatio-temporal features for both inputs via parallel siamese network branches. (ii) adaptive combination through a context-aware fusion gating mechanism that generates dynamic fusion weights to mitigate the imbalanced type effects.
3.3.1. Siamese ST Encoders
To effectively process the distinct characteristics of past- and future-conditioned semantic representations, we adopt a siamese architecture. This involves two identical spatio-temporal (ST) backbone encoders. Each encoder , which can be instantiated with any existing spatio-temporal neural network (Bai et al., 2020; Guo et al., 2019; Wu et al., 2019; Li et al., 2017; Yu et al., 2017; Satorras et al., 2022; Gao and Ribeiro, 2022), is tasked with capturing the complex spatial correlations and temporal dependencies within its respective input stream. The parallel processing yields two specialized representations:
where and encapsulate the high-level spatio-temporal features derived from past and future exogenous contexts, respectively, preparing them for adaptive fusion.
3.3.2. Context-Aware Balancer
Simple fusion or weighted representations will not be able to adapt to dynamic changes in context. Therefore, the representations and are then integrated via a context-aware fusion gate that dynamically computes balancing weights. This gate learns to adjust the contribution of each branch based on the specific input instance, thereby mitigating the imbalance between noise and uncertainty. First, the two representations are element-wise add to form a unified context tensor . An instance-specific gating signal is then generated through a multi-step transformation:
where first aggregates spatial information to create a channel-wise descriptor. This descriptor is then passed through a bottleneck structure composed of two layers, with a ReLU non-linearity, and finally a activation to produce the balancing weight . The final prediction is obtained by a weighted combination of the two branches, where is broadcast appropriately:
In addition, we also use the residual connection technique to supplement the context tensor to promote information propagation. This adaptive gating mechanism allows the model to dynamically prioritize past versus future information based on the input context, effectively balancing their distinct type effects and producing a robust, fused result for prediction.
4. Experiments
In this section, we conduct extensive experiments to investigate the following research questions:
RQ1: Can ExoST outperform exogenous variable modeling methods in accuracy across various task scenarios? (Effectiveness)
RQ2: Can ExoST achieve consistent improvements on various types of existing spatio-temporal backbones? (Universality)
RQ3: Can ExoST maintain its performance in scenarios where exogenous variables are missing or noisy? (Robustness)
RQ4: How does ExoST work? Which components or strategies are crucial? Are these components or strategies sensitive to different desgin or hyper-parameters? (Mechanism)
RQ5: How does ExoST compare in efficiency and parameter count to other methods? (Efficiency & Lightweight)
4.1. Experimental Setup
Datasets. We conducted various experiments on two public real-world spatio-temporal datasets 111https://zenodo.org/records/7308425, which contains hourly measurements collected from monitoring stations in Madrid during the periods 1 January to 30 June 2019 and 1 January to 30 June 2022. Each dataset includes monitoring data from air quality, meteorological, and traffic systems. We consider four forecasting tasks: (i) AQI-19 and AQI-22: predict concentrations in air quality system using 2019 / 2022 data, with traffic and meteorological indicators as exogenous variables; (ii) Speed-19 and Intensity-22: predict traffic speed / intensity in the traffic system using 2019 / 2022 data, with other system indicators as exogenous variables. All dataset are chronologically split into training, validation, and test sets in a 7:2:1 ratio. In addition, it is worth noting that we also include the coding of dates as both past and future exogenous variables as input. More details on dataset, please refer to Appendix A.1.
Baseline. For performance comparison, we categorize existing baselines into three groups: (i) time series models with exogenous variable modeling : TimeXer (Wang et al., 2024), TiDE (Das et al., 2023), and NBEATSx (Olivares et al., 2023); (ii) spatio-temporal models with exogenous variable integration: MAGCRN (Giganti et al., 2024); (iii) general exogenous variable modeling frameworks: ChronosX (Arango et al., 2025), which allows for the substitution of any backbone. Additionally, to evaluate ExoST’s universality, we extensively utilize various classic spatio-temporal backbones, including AGCRN (Bai et al., 2020), GWNet (Wu et al., 2019), GGNN (Satorras et al., 2022), GRUGCN (Gao and Ribeiro, 2022), STGCN (Yu et al., 2017), and DCRNN (Li et al., 2017). More details on methods, see Appendix A.2.
Protocol. By default, we use the previous 24 steps of historical data to predict data for the next 24 steps (i.e., 1-day), and evaluate using mean absolute error (MAE), root mean square error (RMSE), mean relative error (MRE), and mean absolute percentage error (MAPE). Moreover, to measure the long-term generalization of the model in real-world deployments, we also perform rolling evaluations over 2 and 3 days without retraining. For the baselines, we use the parameter settings recommended in their papers and fine-tune them. For ExoST, we set the number of gated experts to 4 and the hidden size to 64. To ensure fairness, each experiment is repeated five times, and results are reported as the mean standard deviation. Best results are highlighted in bold pink and the second-best in underlined blue. For more details on the protocol, see Appendix A.3.
4.2. Effectiveness Study (RQ1)
Table LABEL:tab:rq1_overall_performance shows the performance comparison of our model with various state-of-the-art methods for time series forecasting or spatio-temporal forecasting that utilize exogenous variable modeling. It is worth noting that the first three time series forecasting methods have coupled model designs, while MAGCRN, ChronosX, and our method all use classical AGCRN(Bai et al., 2020) as the backbone network to ensure a fair comparison. Our observations are as follows: ❶ Time series forecasting methods generally achieve the worst results, which is consistent with common spatio-temporal forecasting benchmarks (Shao et al., 2024). This is because they generally do not consider spatial dependencies, resulting in inaccurate predictions. ❷ The ChronosX method, which utilizes a spatio-temporal backbone, despite achieving relative progress, often encounters non-convergence issues on some tasks during training (i.e., Speed-19 and Intensity-22). Despite repeated parameter tuning, this issue cannot be avoided. We attribute this to the design of the underlying Chronos-type (Ansari et al., [n. d.]) time series foundation model. ❸ MAGCRN often achieves suboptimal results because it is specifically designed for spatio-temporal prediction scenarios and still ignores the inconsistent and unbalanced effects of exogenous variables. ❹ Our ExoST achieves competitive performance across all metrics on the four tasks. By effectively selecting and balancing exogenous information, the ExoST can adapt to a variety of input sources, maintaining stable performance.
4.3. Universality Study (RQ2)
We further aim to demonstrate the universality of our ExoST’s “select, then balance” paradigm across various spatio-temporal backbones. Table LABEL:tab:rq1_3day illustrates the performance improvements of our proposed ExoST for 3-day predictions across four tasks and six models. The ✗ column reflects the the performance metrics of the backbone model alone on the target system, while the ✓ column displays the improved results achieved by our ExoST through exogenous variable modeling. The relative improvement is quantified in the column. Based on these results, we make the following observations: ❶ Our ExoST exhibits significant performance improvements across different architectures and datasets, attributed to an effective exogenous variable modeling paradigm. ❷ From the model perspective, the extent of performance improvement varies among different backbone networks, with AGCRN yielding the optimal overall enhancement; however, even the least performing Gwnet model achieved substantial performance gains. ❸ From the task perspective, air quality prediction consistently showed stable performance improvements, as exogenous information like traffic and meteorology significantly impacts it; conversely, traffic speed prediction saw only moderate improvements due to the limited influence of exogenous information such as air quality. We further supplement results for 1- and 2-day predictions in Appendix B.1 and additionally find that ❹ as the prediction horizon extends, our ExoST achieves more significant gains, which also indicates the strong generalization capability of our method.
4.4. Robustness Study (RQ3)
In complex real-world scenarios, exogenous variable data in spatio-temporal systems may not always be collected reliably (for example, signal loss caused by sensor anomalies or maintenance), which can affect prediction accuracy and model robustness. To evaluate the robustness of the ExoST under such adverse conditions, we designed a systematic experiment to simulate real-world data corruption scenarios. Specifically, we randomly replaced exogenous variables with varying degrees of degradation (20%, 40%, 60%, and 80%) with zero values or randomly generated values from a standard normal distribution to evaluate the model’s ability to maintain predictive performance despite exogenous information corruption. The 3-day prediction experiment results shown in Table LABEL:tab:mask_3day reveal several interesting phenomena: ❶ As the proportion of missing data increases, model performance gradually decreases in most cases. However, the overall absolute performance drop is not significant, indicating that our framework can still maintain reasonable performance with limited auxiliary information from exogenous variables. ❷ For some tasks, appropriate missing signals can lead to improved performance, primarily because random data corruption can be viewed as a simple data augmentation strategy. By introducing different data perturbations, the model is forced to learn to predict under a variety of incomplete or noisy data conditions, thereby enhancing its generalization and robustness. We also provide a performance comparison for the 1-day and 2-day scenarios in Appendix B.2, observing similar results.
4.5. Mechanism Study (RQ4)
4.5.1. Ablation Study
We first conducted an ablation study to systematically evaluate the impact of different components on ExoST from two perspectives: data input and model component.
From the data perspective, we aimed to investigate the auxiliary role of different types of exogenous variables in the final prediction. Specifically, due to the universality of date information, here, we treated it as a separate category. As shown in Table LABEL:tab:performance_1day, we observed that ❶ future exogenous variables outperformed past, indicating that forward-looking information provides more direct predictive signals. However, combining both past and future yielded better performance than using either type alone, which suggests that the combination can provide complementary patterns. ❷ We also found that date exogenous variables performed poorly when used in isolation but led to improvements when combined with any other type. ❸ Best performance was achieved by combining all exogenous variables. This demonstrates that ExoST can effectively exploit different information, extracting and balancing contributions from different temporal perspectives and contextual information.
From the model perspective, we aimed to investigate the auxiliary role of different module components in the final prediction. We constructed two variants by removing the MoE Selector and the Context-aware Balancer (achieved through direct element-wise fusion), respectively. As Figure 9 shows, ❶ w/ Selector variant significantly degrades performance in all metrics, demonstrating the critical importance of addressing inconsistent variable effects. Performance degradation indicates that without proper semantic alignment and adaptive selection mechanisms, the model struggles to disentangle heterogeneous exogenous variables from endogenous variables, leading to feature space competition and channel entanglement issues. ❷ w/ Balancer variant causes even more pronounced performance degradation than w/ Selector variant, highlighting that temporal distribution differences between past and future exogenous information pose a more fundamental challenge, requiring sophisticated balancing mechanisms to handle noise-contaminated past information and uncertainty-amplified future forecasts.
More experimental results are shown in Appendix B.3.1.
4.5.2. Parameter Study.
We then conducted a parameter study to evaluate the impact of different parameter on ExoST. As Figure 5 shows, we can observe that ❶ The hidden size of 64 achieves optimal performance in all prediction horizons, with larger sizes leading to performance degradation due to overfitting. ❷ Expert number of 4 provides the best balance between model capacity and computational efficiency. Too few experts (one or two) fail to capture the diverse semantic patterns needed for disentangling inconsistent variable effects, while excessive experts (eight) lead to overselection where redundant experts compete for similar patterns, diminishing the effectiveness of the mixture-of-experts gating mechanism designed to mitigate channel entanglement. These observations validate the effectiveness of our lightweight design choices in the latent-space gated expert module.
4.5.3. Strategy Study.
We also use several alternative strategies to test the rationality of our design. These include: (i) Shared Parameter: Two siamese branches share parameters to learn a unified spatio-temporal representation. (ii) Simple Weight: Adopts a predefined balancing weight () to fuse the outputs from two siamese branches. (iii) Learnable Weight: Dynamically adjusts the relative importance of past and future information through learnable parameters. (iv) Attention: Utilizes cross-attention mechanisms to dynamically balance the contributions from two siamese branches. As shown in Figure 6, we observe that: ❶ ExoST achieves competitive performance in other strategies, demonstrating the effectiveness of our design choice. ❷ Simple Weight shows significant performance degradation due to its static balancing strategy that does not adapt to varying input contexts. ❸ Attention and Learnable Weight strategies underperform while increasing computational complexity, as attention mechanisms introduce excessive parameter overhead and learnable weights lack the sophisticated context modeling capabilities needed to handle nuanced differences between exogenous variables of the past and future. More experimental results are shown in Appendix B.3.2.
4.6. Efficiency & Lightweight Study (RQ5)
We compared the differences in training time, memory footprint, and performance between ExoST and other time series or spatio-temporal forecasting models that incorporate exogenous variables on the AQI-19 dataset. Due to varying convergence epochs, for a fair comparison, we only present the average training time for the first 100 epochs. As shown in Figure 7, compared to several other methods, ExoST demonstrates a balance of performance, efficiency, and model parameter count. Furthermore, we can observe that: ❶ compared to time series models with different backbones, despite its lightweight model design, it leads to poor performance. ❷ compared to spatio-temporal method and the general exogenous variables modeling method with the same backbone (i.e., MAGCRN, ChronosX), we achieved further performance improvement while maintaining similar parameters and training time.
5. Conclusion and Future Work
This paper investigates the challenges of modeling exogenous variables for spatio-temporal forecasting and explores effective implementation options. We propose a novel framework, ExoST, which follows a “select, then balance” paradigm. It first dynamically selects and restructures salient signals via a latent space gated expert module to address inconsistent variable effects. It then incorporates a dual-branch architecture with a context-aware weighting mechanism to address imbalanced type effects. Extensive experiments demonstrate its effectiveness, universality, and robustness. In future work, we aim to explore how to broaden the modal availability of exogenous variables and their feasibility in data-scarce scenarios.
Supplementary material
Select, then Balance: A Plug-and-Play Framework for Exogenous-Aware Spatio-Temporal Forecasting
[…]
[sections] Table of Contents
[sections]l1
Appendix A Experimental Details
A.1. Datasets Details
The statistics of the two real-world datasets used in this paper are summarized in Table 1.
| Attribute | Madrid-19 | Madrid-22 | ||||
| Time range | 01/01/2019 – 30/06/2019 | 01/01/2022 – 30/06/2022 | ||||
| Frames | 4344 | 4344 | ||||
| Sampling Rate | 1 hour | 1 hour | ||||
| Data points | 2.03 M | 2.03 M | ||||
| Variables | 20 | 20 | ||||
| Variable Domain |
Graph Construction for GNN-based Models. We classify the GNN models into adaptive graph construction and predefined graph structure based on their graph construction approaches and perform graph construction on datasets accordingly. For adaptive graph construction models including AGCRN, GraphWaveNet, GGNN, and MAGCRN, the adjacency matrix is computed by embedding the learnable node. Specifically, given node embeddings where is the embedding dimension, AGCRN computes the adjacency matrix as , where computes the similarity between all pairs of nodes, ReLU ensures non-negative similarities, and softmax normalizes the weights to create a valid adjacency matrix. This approach allows the model to automatically discover the latent spatial dependencies between the monitoring stations without requiring prior knowledge of their geographical relationships. GWNet uses separate source and target embeddings to compute the adjacency matrix as , allowing more flexible spatial relationship modeling. GGNN employs a fully connected graph structure where for all , enabling dense spatial attention across all pairs of node. MAGCRN inherits the adaptive graph construction capability from AGCRN, automatically learning the graph structure through node embeddings. For predefined graph structure models that include STGCN, DCRNN, and GRUGCN, due to the lack of geographic location, we follow the mainstream practice 222Torch-Spatiotemporal and construct the adjacency matrix based on the Pearson correlation coefficient of the target time series. Given the target variable at time , we compute the correlation matrix where represents the Pearson correlation coefficient between nodes and . The adjacency matrix is then constructed as , where the operation retains the most correlated connections for each node, specifically choosing .
Variable Scales and Units. For the AQI-19 and AQI-22, the target variable ranges from 0 to 328 g/m3 in Madrid-19 and 0 to 625 g/m3 in Madrid-22, with mean values of 36.6 and 28.0 g/m3 respectively. For the Speed-19 and Intensity-22, the target variables have various scales: the average traffic speed in Madrid-19 ranges from 0 to 3.57 with a mean of approximately 1.2 , while the traffic intensity in Madrid-22 ranges from 56.5 to 289.88 vehicles/hour with a mean of approximately 170 vehicles/hour. Traffic variables show substantial scale differences: traffic intensity ranges from to , traffic occupation time varies from 0.5% to 5.76%, traffic load ranges from 6.5% to 28.12%, and traffic average speed ranges from to . Exogenous variables maintain consistent scales across all tasks, with meteorological variables serving as future covariates and traffic variables (excluding the target) serving as past covariates. To handle these diverse scales effectively, we apply StandardScaler normalization to all variables, ensuring that each variable has zero mean and unit variance.
Temporal Variable Encoding. We employ consistent temporal encoding in four forecasting tasks using sinusoidal transformations for periodic features and one-hot encoding for categorical information. For continuous periodic features, given a timestamp , the hour encoding is calculated as and , and the month encoding as and . For categorical temporal information, each weekday is represented as a binary vector , where if the day corresponds to the -th day of the week (Monday=0, Sunday=6), and otherwise. These temporal features are added as universal exogenous variables to four tasks.
A.2. Baseline Details
In this appendix, we provide detailed descriptions of the advanced exogenous variable modeling methods and different spatio-temporal backbone models used in our default evaluation.
A.2.1. Exogenous Variable Modeling Methods
These methods are grouped into three main categories based on their approach to handling exogenous variables.
TS Models w/ Exogenous Variable Integration.
TiDE (Das et al., 2023): TiDE is an encoder–decoder model built on multilayer perceptrons (MLPs) that fuses historical time-series data and covariates via dense MLP residual blocks to map them into low-dimensional representations, then generates forecast sequences by incorporating future covariates. https://github.com/google-research/google-research/tree/master/tide
TimeXer (Wang et al., 2024): TimeXer is a Transformer model tailored for time series forecasting with exogenous variables: it captures temporal dependencies in the endogenous series via patch-level self-attention and effectively integrates exogenous information through variable-level cross attention and learnable global endogenous tokens to produce high-quality predictions of the target series. https://github.com/thuml/TimeXer
NBEATSx (Olivares et al., 2023): NBEATSx extends the original N-BEATS by incorporating convolutional encoding substructures for time-varying and static exogenous variables within a hierarchical stacked residual block architecture, enabling interpretable decomposition of trend, seasonality, and exogenous factors for high-quality target-series forecasting. https://github.com/cchallu/nbeatsx
ST Models w/ Exogenous Variable Integration.
MAGCRN (Giganti et al., 2024): MAGCRN is a spatio-temporal model that extends AGCRN into a dual-path architecture, using separate AGCRN blocks to encode historical observations and future exogenous covariates. https://github.com/polimi-ispl/MAGCRN
General Exogenous Variable Integration Framework.
ChronosX (Arango et al., 2025): ChronosX extends MAGCRN by integrating two exogenous adaptation modules: Instance-wise Interpolation Block (IIB) and Offset Integration Block (OIB). IIB adaptively interpolates exogenous inputs across time steps, while OIB injects time-dependent bias offsets into the output layer, enabling more flexible integration of future covariates. Due to the limited public availability of the code, we implemented the IIB and OIB modules based on the formulas provided in the referenced paper. https://github.com/amazon-science/chronos-forecasting/tree/chronosx
A.2.2. Spatio-Temporal Backbone Models
To ensure fair use of these models, we use a popular spatio-temporal forecasting benchmark library: https://github.com/TorchSpatiotemporal/tsl.
AGCRN (Bai et al., 2020): AGCRN is an adaptive graph-based recurrent model that captures fine-grained spatio-temporal dependencies by jointly learning node-specific patterns and data-driven graph structures, eliminating the need for pre-defined adjacency matrices.
GWNet (Wu et al., 2019): GWNet is a spatial-temporal model that combines diffusion-based graph convolutions with dilated causal temporal convolutions, and incorporates a learnable self-adaptive adjacency matrix to capture latent spatial relationships from data.
GGNN (Satorras et al., 2022): GGNN is a recurrent graph-based model that applies gated update mechanisms to iteratively refine node representations through message passing, enabling effective modeling of long-range structural dependencies.
GRUGCN (Gao and Ribeiro, 2022): GRUGCN is a spatio-temporal neural architecture that integrates graph convolutions for spatial feature extraction and gated recurrent units for sequential modeling, effectively capturing dynamic dependencies in traffic forecasting scenarios.
STGCN (Yu et al., 2017): STGCN is a graph-based temporal forecasting model that integrates graph convolutional layers with temporal convolution modules to jointly learn spatial and temporal dependencies in structured time series data.
DCRNN (Li et al., 2017): DCRNN is a graph-based recurrent forecasting model that captures spatial dependencies via diffusion convolutions on directed graphs and models temporal dynamics through a sequence-to-sequence recurrent architecture.
A.3. Protocol Details
A.3.1. Metrics Detail.
We use different metrics such as MAE, RMSE, MRE, and MAPE. Formally, these metrics are formulated as follows:
| MAE | RMSE | ||||
| MRE | MAPE |
where represents the indices of all observed samples, denotes the -th actual sample and is the corresponding prediction. The mean relative error is defined as the MAE divided by the norm of the target window.
A.3.2. Parameter Detail.
The model is trained for 500 epochs using the AdamW optimizer with a cosine annealing learning rate scheduler. The learning rate is initialized at and decays to over the training period. A batch size of 512 is used, incorporating mixed-precision arithmetic and Z-score normalization for target variables. Early stopping is implemented to prevent overfitting, halting training if the validation MAE does not improve for 30 consecutive epochs. All experiments are carried out on a Linux server equipped with a 1 AMD EPYC 7763 128-Core Processor CPU with 256GB memory and 4 NVIDIA RTX A6000 GPUs, each with 48GB memory.
Appendix B More Results
B.1. More Result on Universality Study
| Model | AGCRN | GWNet | GGNN | GRUGCN | STGCN | DCRNN | |||||||||||||
| w/ ExoST | ✗ | ✓ | (%) | ✗ | ✓ | (%) | ✗ | ✓ | (%) | ✗ | ✓ | (%) | ✗ | ✓ | (%) | ✗ | ✓ | (%) | |
[…]
| Model | AGCRN | GWNet | GGNN | GRUGCN | STGCN | DCRNN | |||||||||||||
| w/ ExoST | ✗ | ✓ | (%) | ✗ | ✓ | (%) | ✗ | ✓ | (%) | ✗ | ✓ | (%) | ✗ | ✓ | (%) | ✗ | ✓ | (%) | |
[…]
We provide more experimental results on universality (i.e., for 1-day and 2-day scenarios), as shown in Tables 2 and 3. We observe that while the ExoST exhibits some negative gains in the 1-day scenario, it exhibits improved performance as the prediction length increases, demonstrating the superior generalization of our method. The remaining observations show similar analytical results to those in the main text.
B.2. More Result on Robustness Study
| Datasets | AQI-19 | Speed-19 | Intensity-22 | ||||||||||
| Masking Strategy | MAE | RMSE | MAPE(%) | MRE(%) | MAE | RMSE | MAPE(%) | MRE(%) | MAE | RMSE | MAPE(%) | MRE(%) | |
| Zero | 20%-Mask | ||||||||||||
| 40%-Mask | |||||||||||||
| 60%-Mask | |||||||||||||
| 80%-Mask | |||||||||||||
| Random | 20%-Mask | ||||||||||||
| 40%-Mask | |||||||||||||
| 60%-Mask | |||||||||||||
| 80%-Mask | |||||||||||||
| No Masking | |||||||||||||
| Datasets | AQI-19 | Speed-19 | Intensity-22 | ||||||||||
| Masking Strategy | MAE | RMSE | MAPE(%) | MRE(%) | MAE | RMSE | MAPE(%) | MRE(%) | MAE | RMSE | MAPE(%) | MRE(%) | |
| Zero | 20%-Mask | ||||||||||||
| 40%-Mask | |||||||||||||
| 60%-Mask | |||||||||||||
| 80%-Mask | |||||||||||||
| Random | 20%-Mask | ||||||||||||
| 40%-Mask | |||||||||||||
| 60%-Mask | |||||||||||||
| 80%-Mask | |||||||||||||
| No Masking | |||||||||||||
B.3. More Result on Mechanism Study
B.3.1. Ablation Study.
| P | F | D | MAE | RMSE | MAPE (%) | MRE (%) |
[…]
| P | F | U | MAE | RMSE | MAPE (%) | MRE (%) |
| ✓ | ✓ | - | 43.71±0.70 | |||
[…]
From the data perspective, we also evaluated the impact of exogenous variables on ExoST’s performance for 2-day and 3-day forecasting on the AQI-19 , as shown in Tables 6 and Tables 7. We have further observation: ❶ Consistent with the 1-day findings, future covariates outperform past covariates, while data variables alone produce limited accuracy but enhance performance when combined with others, with the best results achieved by integrating all variables, demonstrating ExoST’s ability to synthesize diverse temporal information. ❸ In 2-day forecasting, combining past and future covariates significantly improves percentage-based metrics, indicating a balanced contribution from historical and forward-looking signals. ❷ For 3-day forecasting, combining future and data covariates almost matches the full model, highlighting the significant role of the immediate temporal context in short-term predictions. ❹ Across 1-day to 3-day horizons, the most critical trend is increasing the importance of the past covariates as the forecast horizon extends, complementing the dominant predictive power of future covariates.
B.3.2. Strategy Study.
We first formalize the four strategies in detail through formulas:
Sharing Parameter: A single spatio-temporal encoder is re-used for both exogenous streams:
The forecast is obtained by averaging the two outputs:
Simple Weight: Each stream is processed by its own encoder 3.2. The final prediction employs fixed, equal weights:
Learnable Weight: Let be a learnable vector, normalized by softmax:
Given the two outputs of the backbone, the fusion is
where the last term is a residual connection.
Attention: Let denote the intermediate hidden states produced by the two branches before the final read-out layer. For each directional attention we use shared projection matrices .
Future Past
Past Future
Final Prediction
As Figures 10 and 11 show, we can observe that: ❶ In the 2-day horizon, ExoST maintains its advantage, while Simple Weight and Learnable Weight exhibit amplified weaknesses: the former’s static weighting fails to handle increased uncertainty in longer forecasts, and the latter, despite its adaptable weights, amplifies noise from future exogenous signals due to lacking day-to-day context modeling. ❷ Althougn all methods show increasing errors across horizons of 1-day to 3-day, ExoST consistently suppresses error accumulation better than alternatives.
Appendix C More Disscuison
C.1. Limitation
In this paper, we propose a new framework for modeling exogenous variables in spatio-temporal forecasting: ExoST. This “select, then balance” design addresses the inconsistent effects of variables and the imbalanced effects of types. Extensive experiments demonstrate the superiority of our method in terms of effectiveness, universality, and robustness. Although we have made some attempts in this direction, there are still some limitations that need to be addressed:
First, our method assumes that spatio-temporal exogenous variables are available in real-time and are synchronized with the target system. In reality, it may not be possible to acquire real-time exogenous information at all times. Even when available, differing equipment may lead to inconsistent and mismatched sampling rates. While we conducted preliminary experiments using random masking to simulate asynchronous and non-real-time data, and our model showed some robustness, a deeper investigation into this issue is left for future work.
Second, our model currently assumes that spatio-temporal exogenous variables are numerical sensor signals. However, real-world scenarios often involve diverse, heterogeneous exogenous information, such as language and visual information. Recent research has explored using Large Language Models (LLMs) for time series (jin, 2024; Pan et al., 2024; Sun et al., [n. d.]; Jia et al., 2024; Chang et al., 2025; Liu et al., 2025b) and spatio-temporal forecasting (Liu et al., 2025a, 2024; Li et al., 2024; Xu and Liu, 2025), hoping that LLMs can gain numerical prediction capabilities through various alignment techniques. However, some studies have raised significant doubts about these approaches (Tan et al., 2024). Different from these views, we believe that language can be more effectively viewed as sparse, discrete exogenous auxiliary information. The key challenge lies in how to leverage this information to enhance various spatio-temporal tasks.
C.2. Future Direction
Based on these analyses, we identify two key areas for future work:
For spatio-temporal forecasting tasks, we believe the core challenge is to use the semantic representations of exogenous variables to provide an auxiliary spatio-temporal backbone network with interventional information, allowing it to respond to dynamic changes. This approach is distinct from using language models as the primary prediction backbone.
For spatio-temporal reasoning tasks, we believe the core essence is to activate the intrinsic capabilities of language models to analyze the underlying causes of spatio-temporal changes and provide logical knowledge-based answers.

### v2

@@
-Select, then Balance: A Plug-and-Play Framework for Exogenous-Aware Spatio-Temporal Forecasting
-Abstract.
-Spatio-temporal forecasting aims to predict the future state of dynamic systems and plays an important role in multiple fields.
-However, existing solutions only focus on modeling using a limited number of observed target variables.
-In real-world scenarios, exogenous variables can be integrated into the model as additional input features and associated with the target signal to promote forecast accuracy.
-Although promising, this still encounters two challenges: the inconsistent effects of different exogenous variables to the target system, and the imbalance effects between historical variables and future variables.
-To address these challenges, this paper introduces ExoST, a novel framework for modeling exogenous variables in spatio-temporal forecasting, which follows a “select, then balance” paradigm.
-Specifically, we first construct a latent space gated expert module, where fused exogenous information is projected into a latent space to dynamically select and recompose salient signals via specialized sub-experts.
-Furthermore, we design a siamese network architecture in which recomposed representations of past and future exogenous variables are fed into dual-branch spatio-temporal backbones to capture dynamic patterns.
-The outputs are integrated through a context-aware weighting mechanism to achieve dynamic balance during the modeling process.
-Extensive experiments on real-world datasets demonstrate the effectiveness, generality, robustness, and efficiency of our proposed framework.
-1. Introduction
-Spatio-temporal forecasting aims to model the dynamic evolution of spatial entities over time by leveraging historical spatio-temporal sequences to predict future states over multiple time steps (Fang et al., 2025; Shao et al., 2023; Chen and Liang, 2025a, b).
-This task finds critical applications in domains such as traffic management (Xia et al., 2025; Avila and Mezić, 2020), environmental monitoring (Liang et al., 2023; Wang et al., 2020), and weather forecasting (Yang et al., 2024; Bi et al., 2023).
-In recent years, researchers have developed various spatio-temporal (ST) neural encoders (Jin et al., 2023b, a) by combining building blocks with different relational inductive biases, such as graph convolution operator (Ma et al., 2025; Yu et al., 2017; Li et al., 2017) for spatial dependencies and sequence operator (Wu et al., 2019; Guo et al., 2019) for temporal patterns.
-However, these approaches predominantly rely on modeling limited observations of target system variables, focusing mainly on encoding the spatial and temporal dynamics of target signals, while overlooking a widely available yet underutilized source of information in real-world forecasting systems: spatio-temporal exogenous variables.
-Spatio-temporal exogenous variables refer to contextual information that is related to the prediction system but originates from external dynamic systems (Berkani et al., 2023).
-Such variables may include historical information that co-occur with the target system as well as prior information about future time points provided by external auxiliary forecasting systems.
-These factors serve as critical complementary signals in prediction tasks, helping to reduce uncertainty while offering interpretability (Huang et al., 2025; Das et al., 2023; Wang et al., 2024; Arango et al., 2025; Olivares et al., 2023; Giganti et al., 2024).
-As illustrated in Fig.
-1 (a), consider the sensor network of air quality monitoring stations in New York City.
-For any given monitoring station, air quality forecasting typically focuses on modeling various indicators of pollutant emissions (e.g., NO2, CO2) to predict future air conditions.
-However, air quality at a particular location may also be significantly influenced by surrounding contextual information.
-For instance, historical traffic congestion data in the region (e.g., traffic flow) could lead to sudden spikes in pollutant levels, while stable weather conditions predicted for the future (e.g., wind speed) may indicate a risk of continued deterioration in air quality.
-Despite the considerable potential of incorporating exogenous information, two major challenges persist in practice:
-✦ Challenge I: Inconsistent Variable Effects: In real-world scenarios, exogenous variables often exert varying degrees of influence on the prediction target.
-As shown in Figure 1(b), different exogenous variables display distinct temporal dynamics and correlations with the target signal.
-Some variables (e.g., traffic flow) may have substantial predictive value, while others (e.g., traffic incidents) may be marginally relevant or even introduce noise.
-Fig.
-2 (a) further illustrates this phenomenon by showing the Pearson correlation coefficients between various exogenous variables and the target variable across randomly selected stations in the AQI-19 dataset.
-Despite this variability, existing approaches often adopt a one-size-fits-all strategy (Arango et al., 2025; Giganti et al., 2024), directly concatenating exogenous variables with input features, without selectively modeling their individual contributions.
-This lack of mechanism-aware integration limits performance and can even degrade predictive accuracy.
-✦ Challenge II: Imbalanced Type Effects: In addition to inter-variable disparities, there also exist imbalanced type differences between past and future exogenous variables from different temporal contexts.
-As shown in Fig.
-1(c), exogenous information may stem from historically perceived signals (e.g., past traffic data) or from proactively forecasted future signals (e.g., weather predictions), leading to temporal distribution asymmetry and context-dependent impacts on prediction outcomes.
-Fig.
-2 (b) clearly demonstrates the distinct distribution patterns of normalized values across different types of exogenous variables in the AQI-19 dataset.
-Existing methods typically adopt joint modeling strategies that fail to distinguish the distinct influences of these two types of exogenous variables on the target system (Wang et al., 2024; Das et al., 2023; Olivares et al., 2023; Zhou et al., 2025).
-Moreover, even under decoupled modeling settings, simple weighted fusion is insufficient to effectively balance their heterogeneous effects.
-To address the aforementioned challenges, we propose a novel exogenous variable-enhanced spatio-temporal forecasting framework, ExoST, which is built upon a core “select, then balance” paradigam to systematically exploit and integrate the contributions of exogenous variables, thereby improving overall predictive performance.
-The framework adopts a modular and flexible design that allows seamless integration with a wide range of mainstream spatio-temporal backbone networks.
-Specifically, ❶ we introduce an adaptive selection mechanism through a latent-space gated expert module, which projects both endogenous and multi-source exogenous variables into a condition latent space and activates specialized expert sub-networks via a gating mechanism.
-This enables the model to automatically select and combine the most informative exogenous factors.
-The module exhibits strong representational capacity and adaptability, effectively capturing relevant signals while suppressing irrelevant or redundant information, thus alleviating inconsistencies caused by variable heterogeneity.
-To address the imbalance between types of exogenous variables, ❷ we design a dual-branch siamese network architecture that independently models past and future exogenous information.
-These recomposed representations are fed into dual versions of any given spatio-temporal backbone to learn evolutionary patterns, and a context-aware balancer is employed to dynamic fusion and produce the final prediction.
+Select, then Balance: Exploring Exogenous Variable Modeling of Spatio-Temporal Forecasting
+Abstract
+Spatio-temporal (ST) forecasting is critical for dynamic systems, yet existing methods predominantly rely on modeling a limited set of observed target variables.
+In this paper, we present the first systematic exploration of exogenous variable modeling for ST forecasting, a topic long overlooked in this field.
+We identify two core challenges in integrating exogenous variables: the inconsistent effects of distinct variables on the target system and the imbalance effects between historical and future data.
+To address these, we propose ExoST, a simple yet effective exogenous variable modeling general framework highly compatible with existing ST backbones that follows a “select, then balance” paradigm.
+Specifically, we design a latent space gated expert module to dynamically select and recompose salient signals from fused exogenous information.
+Furthermore, a siamese dual-branch backbone architecture captures dynamic patterns from the recomposed past and future representations, integrating them via a context-aware weighting mechanism to ensure dynamic balance.
+Extensive experiments on real-world datasets demonstrate the ExoST ’s effectiveness, universality, robustness, and efficiency.
+onedeanxxx@gmail.com, ∗yuxliang@outlook.com
+1 Introduction
+Spatio-temporal forecasting aims to learn the evolution of structured systems whose states vary across both space and time (shao2024exploring; chen2025eac; chen2025learning).
+This task is fundamental to traffic management (avila2020data), environmental monitoring (liang2023airformer), and meteorology (bi2023accurate).
+Recent advances have focused on spatio-temporal (ST) encoders (jin2023spatio; jin2024survey) that integrate graph neural operators (yu2017spatio; ma2025less) for learning spatial dependencies and sequence operators (wu2019graph; guo2019attention; gao2023spatio; han2024bigst) for modeling temporal patterns.
+However, existing methods predominantly focus on the dynamics of target variables, largely neglecting a critical source of auxiliary information: spatio-temporal exogenous variables.
+Spatio-temporal exogenous variables encompass contextual data from external dynamic systems (berkani2023spatio), including historical covariates and auxiliary future forecasts.
+These signals complement target predictions by reducing uncertainty and enhancing interpretability (olivares2023neural; wang2024timexer; huang2025exploiting; Arango2025ChronosX).
+As illustrated in Figure 1 (a), air quality forecasting relies not only on historical pollutant indicators (e.g., NO2, CO2) but also on the environmental context.
+For instance, historical traffic congestion may explain sudden emission spikes, while future weather forecasts (e.g., wind speed) serve as precursors for pollutant dispersion or accumulation.
+Despite the promise of exogenous information, their systematic utilization in ST forecasting remains surprisingly overlooked. We attribute this gap to two primary challenges:
+Challenge I: Inconsistent Variable Effects: Exogenous variables exhibit heterogeneous correlations with prediction targets.
+As shown in Figure 1 (b) (Pearson correlation coefficients between exogenous variables and the target variable across randomly stations in AQI-19), variables vary from highly predictive (e.g., traffic flow) to irrelevant or noisy (e.g., traffic incidents).
+Existing approaches typically employ a uniform concatenation strategy (giganti2024back; Arango2025ChronosX), which fails to distinguish informative signals from noise.
+This lack of selective modeling limits representation capability and degrades predictive accuracy.
+Challenge II: Imbalanced Type Effects: In addition to inter-variable disparities, distinct distributional asymmetry exists between historical signals and proactive future forecasts.
+Figure 1 (c) highlight the disparate distribution patterns of these data types in AQI-19.
+Current methods often rely on joint modeling or simple weighted fusion (das2023long; olivares2023neural; wang2024timexer; zhou2025crosslinear), which fail to effectively decouple and balance the different impacts of past and future contexts on the target system.
+To address these limitations, we propose ExoST, a novel framework governed by a “select-then-balance” paradigm designed to systematically exploit exogenous variables.
+ExoST features a modular design compatible with diverse spatio-temporal backbones.
+We introduce a Latent Space Gated Expert Module that projects endogenous and exogenous variables into a shared latent space, utilizing a gating mechanism to adaptively select informative factors while suppressing noise.
+To mitigate type imbalance, we design a dual-branch Siamese architecture that decouples the modeling of past and future exogenous contexts.
+These representations are processed by dual backbones and integrated via a context-aware balancer to generate the final prediction.
 In summary, our key contributions are as follows:
-Adaptive Selection Mechanism for Inconsistent Variable Effects: We propose a Latent Space Gated Expert Module that effectively handle the latent associations among multi-source exogenous variables.
-Through an expert selection mechanism, it identifies and emphasizes the most predictive external signals, enhancing the quality of the semantic representation input.
-Balance Modeling Strategy for Imbalanced Type Effects: We design a Siamese network architecture that separately models past and future exogenous information branch, enabling dynamic prediction balancing via a context-aware fusion gating mechanism.
-A Flexible Plug-and-Play Framework: ExoST provides a general and flexible enhancement framework for spatio-temporal forecasting.
-It can be broadly applied to various backbones and consistently outperforms existing methods across multiple benchmarks, demonstrating both practical value and scalability.
-2. Preliminaries
-2.1. Related Work
-2.1.1. Spatio-Temporal Forecasting.
+Problem Level: We present the first systematic formalization and exploration of exogenous variable modeling for ST forecasting, addressing a long-neglected gap therein.
+Technical Level: We identify the key challenges in modeling exogenous variables and propose ExoST, a simple yet effective general framework with targeted modules.
+ExoST demonstrates high compatibility with existing spatio-temporal backbones without any modification.
+Evaluation Level: We systematically evaluate ExoST across multiple aspects, verifying its efficacy, universality, robustness, design rationality, efficiency, and lightweight features.
+2 Related Work
+Spatio-Temporal Forecasting.
 Predicting future spatio-temporal signals from historical observations, a spatially extension of multivariate time series forecasting, has evolved significantly.
-Early statistical (Box and Pierce, 1970; Biller and Nelson, 2003) or independent time series models (Nie et al., 2023; Wu et al., 2022; Qiu et al., 2025; Seo and Lim, 2025; Fan et al., 2025) at each location proved insufficient for complex spatio-temporal dynamics (Shao et al., 2024), which involve intricate spatial correlations and heterogeneity (Chen and Liang, 2025a, b).
-This led to hybrid architectures (Cao et al., 2025; Kong et al., 2024; Ma et al., 2025; Liu et al., 2023; Shao et al., 2022; Wu et al., 2019; Guo et al., 2019; Li et al., 2017) that integrate spatial (e.g., graph operators for node interactions) and temporal (e.g., sequence operators for long-term trends) modeling to jointly learn rich spatio-temporal representations for accurate prediction.
-However, a critical limitation persists: these models largely focus on the system’s internal dynamics while neglecting exogenous variables—external inputs that are independent of the system’s state but directly influence its evolution, and whose inclusion is vital for robust and accurate predictions.
-2.1.2. Exogenous Variables Modeling.
-In practical time series forecasting scenarios, exogenous information is already integrated into the model as additional input features and concatenated with the target signal, enabling the model to learn the correlation between external factors and the target variable.
-Earlier studies explored statistical methods such as ARIMAX (Williams, 2001) and SARIMAX (Vagropoulos et al., 2016) that can understand the relationship between exogenous and endogenous sequences and autoregression.
-Advanced deep learning models discuss how to use future values of exogenous variables (Das et al., 2023; Olivares et al., 2023) and how to adapt to time lags and missing exogenous variable data scenarios (Wang et al., 2024; Zhou et al., 2025).
-In addition, some works (Huang et al., 2025; Liu et al., 2024; Jin et al., [n.
-d.]) also discuss how to use language models to align the semantic information of exogenous variables for enhancement.
-Limited works have explored the integration of exogenous variables in spatio-temporal forecasting scenarios, but they only focus on specific task scenarios and model structure design (Giganti et al., 2024; Dong et al., 2023; Zagouras et al., 2015; Lindström et al., 2014).
-In contrast, we systematically explore how to effectively model exogenous variables in spatio-temporal forecasting scenarios and adaptively weigh the impact of different exogenous variables.
-2.2. Problem Formulation
+Early statistical (box1970distribution; biller2003modeling) or independent time series models (nie2023time; wu2023timesnet; qiu2025duet) at each location proved insufficient for complex spatio-temporal dynamics (shao2024exploring), which involve intricate spatial correlations and heterogeneity (chen2025eac; chen2025learning).
+This led to hybrid architectures (li2017diffusion; wu2019graph; liu2023spatio; chen2025information; cao2025spatiotemporal; ma2025robust) that integrate spatial (e.g., graph operators for node interactions) and temporal (e.g., sequence operators for long-term trends) modeling to jointly learn rich spatio-temporal representations for accurate prediction.
+However, a critical limitation persists: these models largely focus on the system’s internal dynamics while neglecting exogenous variables, external inputs that are independent of the system’s state but directly influence its evolution, and whose inclusion is vital for robust and accurate predictions.
+Exogenous Variables Modeling.
+Distinct from our ST forecasting setting, which models multiple systems as exogenous variables, existing literature primarily focuses on simple covariate modeling within multivariate time series forecasting to capture correlations between observations and target variables.
+Early statistical methods, such as ARIMAX (williams2001multivariate) and SARIMAX (vagropoulos2016comparison), addressed the relationships among exogenous series, endogenous series, and autoregressive components.
+More recent deep learning approaches have investigated utilizing future exogenous values (das2023long; olivares2023neural), handling time lags and missing data (wang2024timexer; zhou2025crosslinear), and leveraging language models to align semantic information for representation enhancement (liu2024spatial; huang2025exploiting).
+While limited research integrates exogenous variables into ST forecasting, as summarized in Table 1, these efforts are restricted to specific ST tasks and model architectural designs (lindstrom2014flexible; zagouras2015role; dong2023spatiotemporal; giganti2024back; ma2025causal; ma2025spatiotemporal; wang2025unifying).
+In contrast, we provide the first systematic formalization and exploration of modeling exogenous variables within this context, proposing a simple yet effective general framework.
+| Category | Example Works | Typical Integration | Universality | Simplicity | Robustness | Efficiency |
+| Statistical Models | ARIMAX (williams2001multivariate), SARIMAX (vagropoulos2016comparison) | Linear Regression | ✗ (Strict Assump.) | ✓ | ✗ | |
+| Specific TS Models | TiDE (das2023long), TimeXer (wang2024timexer) | Concatenation / Cross-Attn | ✗ (TS specific) | ✗ | ✓ | |
+| LLM-based Models | ST-LLM (liu2024spatial), ExoLLM (huang2025exploiting) | Semantic Embedding | ✓ (Semantic Bridge) | ✗ | ✗ | |
+| Specific ST Models | SEEformer (dong2023spatiotemporal), MAGCRN (giganti2024back) | Task-Specific Fusion | ✗ (ST specific) | ✗ | ✓ | |
+| General ST Framework | ExoST (Ours) | Select, then Balance | ✓ (Any ST Backbone) | ✓ | ✓ | ✓ |
+3 Preliminaries
 Definition 1.
 (Spatio-Temporal Data).
-Spatio-temporal data refers to observational records that capture how dynamic systems evolve in space and time.
-Each data point is associated with a specific geographic location and a specific timestamp.
-This type of data can usually be represented as a temporal sequence , where each is a snapshot of the spatial observation at time .
-The spatial is a sensor network consisting of discrete nodes, and each node has features at time (e.g., pollutant indicators CO2, NO2, etc.
-from a dynamic air monitoring systems.).
+Spatio-temporal data records the evolution of dynamic systems across space and time.
+Formally, we represent this as a sequence , where denotes the system state at time .
+The spatial domain is modeled as a network with nodes, where each node observes features (e.g., pollutant indicators such as NO2, etc.
+in air quality system).
 Definition 2.
 (Exogenous Variables).
-Exogenous variables are external system factors that are independent of the spatio-temporal forecasting and are free of the internal dynamics.
-During the modeling process, these variables are fed into the model as known inputs (whether actual observations or predicted values).
-Typically, these data hold the same structure as spatio-temporal data, denoted as , and can be further subdivided into past exogenous variables (e.g., traffic flow and traffic speed at past moments) and future exogenous variables (e.g., future temperature and precipitation forecasts) according to their time attributes.
-Problem 0.
+Exogenous variables are external factors independent of the target system’s internal dynamics, serving as auxiliary inputs.
+Denoted as , these variables share the structural format of ST data.
+comprises past exogenous (e.g., historical traffic flow) and future exogenous variables (e.g., weather forecasts), distinguished by their temporal availability.
+Problem.
 (ST Forecasting with Exogenous Variables).
-Our goal is to learn an optimal prediction model for predicting future spatio-temporal signals from historical observations, optionally using exogenous inputs (and an optional graph structure prior ).
-Let be a training dataset sampled from a spatio-temporal process, where represents the features of nodes for the past time steps, and represents the target for the future time steps.
-Traditional spatio-temporal prediction optimizes the model parameters by minimizing the expected prediction loss:
-where is the prediction model and measures the prediction error. When exogenous input is available, the goal becomes:
-where and denote the exogenous inputs in the past and future, respectively. and are the corresponding feature dimensions.
-3. Methodology
-3.1. Overview
+We aim to predict future spatio-temporal signals given historical observations, optional exogenous inputs, and a graph prior .
+Let denote the training set, where and represent historical features (length ) and prediction targets (length ), respectively.
+Standard forecasting optimizes parameters by minimizing the expected loss:
+where is the model and denotes the error metric. With exogenous inputs , the objective changes to:
+where and correspond to past and future exogenous features with dimensions and .
+In practice, according to (jin2023spatio; jin2024survey; shao2024exploring), the feature to be predicted is usually only the target variable, i.e., in is equal to 1.
+4 Methodology
 The ExoST is built on the “select-then-balance” paradigm, as shown in Figure 2.
-The process consists of two consecutive stages, which are used to address the two core challenges in exogenous-augmented spatio-temporal forecasting mentioned above: (i) inconsistent variable effects and (ii) unbalanced type effects.
-Select stage (§3.2).
-A lightweight Latent-Space Gated Expert module projects heterogeneous exogenous variables into a condition latent manifold, and then activates a Mixture-of-Experts Selector to dynamically re-weight and re-compose the most meaningful signals.
-The output is two refined, exogenous-aware representations and , one for each type.
-Balance stage (§3.3).
-A Siamese dual-branch architecture processes the two representations in parallel through the existing spatio-temporal backbones, and a Context-Aware Balancer adaptively balances their contributions based on instantaneous context.
-The final forecast is produced by a read-out layer on top of the fused hidden state.
+The process consists of two consecutive stages, which are used to address the two core challenges in exogenous-aware ST forecasting mentioned above: (i) inconsistent variable effects and (ii) unbalanced type effects.
 Both stages are fully differentiable and backbone-agnostic: any modern ST encoder can be plugged in without architectural change.
-3.2. Select: Latent-Space Gated Expert Module
+4.1 Select: Latent-Space Gated Expert
 The selection stage aims to address inconsistent variable effects by extract compact, adaptive semantic representations from input endogenous variables and heterogeneous, multi-source exogenous variables, laying the foundation for the subsequent balancing stage.
-This requires a carefully designed two-step process: (i) semantic alignment of diverse exogenous and endogenous variables in a unified semantic space via conditional embeddings; (ii) adaptive selection through a mixture-of-experts selector that reweights and combines existing features to mitigate antagonism caused by inconsistent variable effects.
-3.2.1. Conditional Embedding
-Conditional embeddings specifically address the semantic space alignment of endogenous variables with two types of exogenous variables.
-For each exogenous type (i.e., past, future), we first project the raw exogenous tensor and the target history input into a condition latent space via:
-where is implemented as affine-transformation block:
-where , are learned projections that align heterogeneous feature dimensions into a latent semantic space.
-Specifically, consistent with common settings in existing literature (Shao et al., 2024; Jin et al., 2024), the lengths of past and future steps are typically uniform.
-Therefore, to maintain this setting, we perform a direct element-wise add fusion.
-For more arbitrary cases, zero-padding preprocessing trick can also easily ensure alignment.
-Subsequently, non-linear activation functions project endogenous variables into latent conditional semantic spaces, derived from different types of exogenous variables, yielding conditional representations and , respectively.
-3.2.2. Mixture-of-Experts Selector
-As discussed, inconsistent variable effects between exogenous and endogenous variables can induce antagonistic or synergistic semantic degeneration.
-This manifests as channel entanglement in the derived conditional representations, where semantic mixing may lead to rank-deficient representations, complicating the learning process.
-To address this, we attach a mixture-of-experts gating (Jacobs et al., 1991; Shazeer et al., 2017; Fedus et al., 2022) module on top of each .
-Concretely, we maintain expert projections and a gating network parameterized by a single linear layer followed by function.
-For each conditional representation input, the gating score is applied at different time steps and locations and is calculated as follows:
-where is projection layer projection layer, along with a normalized activation function, calculates each expert’s weight.
-This controls the weighting of each expert for the current input, simulating various semantic combinations.
-The final representation is then obtained through weighted combination:
-This strategy explicitly disentangles representations using a latent bottleneck and selective recombination, rather than forcing the model to learn entangled mappings.
-This latent space gating method allows us to model multiple interactions between exogenous variables and the target representation within the fused conditional space, thereby adaptively acquiring suitable representation input and for the backbone network.
-3.3. Balance: Dual-Branch Adaptive Fusion
-The balancing stage aims to address the unbalanced type effect, that is, how to properly process the two semantic representations in the previous stage.
-This requires appropriately learning spatio-temporal dynamic patterns and effectively fusing them to achieve the final prediction.
-However, a key challenge arises from the inherent difference in temporal distribution between past and future exogenous information, complicating their effective fusion.
-To address this, we designed two crucial steps: (i) independent extraction of dynamic spatio-temporal features for both inputs via parallel siamese network branches.
-(ii) adaptive combination through a context-aware fusion gating mechanism that generates dynamic fusion weights to mitigate the imbalanced type effects.
-3.3.1. Siamese ST Encoders
-To effectively process the distinct characteristics of past- and future-conditioned semantic representations, we adopt a siamese architecture.
-This involves two identical spatio-temporal (ST) backbone encoders.
-Each encoder , which can be instantiated with any existing spatio-temporal neural network (Bai et al., 2020; Guo et al., 2019; Wu et al., 2019; Li et al., 2017; Yu et al., 2017; Satorras et al., 2022; Gao and Ribeiro, 2022), is tasked with capturing the complex spatial correlations and temporal dependencies within its respective input stream.
-The parallel processing yields two specialized representations:
-where and encapsulate the high-level spatio-temporal features derived from past and future exogenous contexts, respectively, preparing them for adaptive fusion.
-3.3.2. Context-Aware Balancer
-Simple fusion or weighted representations will not be able to adapt to dynamic changes in context.
-Therefore, the representations and are then integrated via a context-aware fusion gate that dynamically computes balancing weights.
-This gate learns to adjust the contribution of each branch based on the specific input instance, thereby mitigating the imbalance between noise and uncertainty.
-First, the two representations are element-wise add to form a unified context tensor .
-An instance-specific gating signal is then generated through a multi-step transformation:
-where first aggregates spatial information to create a channel-wise descriptor.
-This descriptor is then passed through a bottleneck structure composed of two layers, with a ReLU non-linearity, and finally a activation to produce the balancing weight .
-The final prediction is obtained by a weighted combination of the two branches, where is broadcast appropriately:
-In addition, we also use the residual connection technique to supplement the context tensor to promote information propagation.
-This adaptive gating mechanism allows the model to dynamically prioritize past versus future information based on the input context, effectively balancing their distinct type effects and producing a robust, fused result for prediction.
-4. Experiments
-In this section, we conduct extensive experiments to investigate the following research questions:
-RQ1: Can ExoST outperform exogenous variable modeling methods in accuracy across various task scenarios? (Effectiveness)
-RQ2: Can ExoST achieve consistent improvements on various types of existing spatio-temporal backbones? (Universality)
-RQ3: Can ExoST maintain its performance in scenarios where exogenous variables are missing or noisy? (Robustness)
-RQ4: How does ExoST work? Which components or strategies are crucial? Are these components or strategies sensitive to different desgin or hyper-parameters? (Mechanism)
-RQ5: How does ExoST compare in efficiency and parameter count to other methods? (Efficiency & Lightweight)
-4.1. Experimental Setup
+It operates in two key stages.
+Conditional Embedding.
+Conditional embeddings specifically address the semantic space and dimension size alignment of endogenous variables with two types of exogenous variables.
+For each type (i.e., past, future), we project the raw exogenous tensor and historical input into a condition latent space:
+where acts as an affine transformation block:
+with learnable projections and .
+Following standard protocols (shao2024exploring; jin2024survey), we assume uniform sequence lengths of past and future steps to enable direct element-wise addition.
+For more arbitrary cases, the zero-padding preprocessing trick can also easily ensure alignment.
+This process yields the conditional representations .
+Gated Expert Selector.
+To mitigate semantic degeneration and channel entanglement arising from inconsistent variable effects, we incorporate a Gated Expert Selector module for each .
+Specifically, we employ expert projections and a gating network .
+The gating scores , which determine the contribution of each expert based on specific spatio-temporal contexts, are computed via:
+where is a learnable projection matrix, along with a normalized activation function, calculates each expert’s weight.
+The refined representation is subsequently synthesized through a weighted aggregation:
+This mechanism explicitly disentangles representations via a latent bottleneck and selective recombination, adaptively deriving robust inputs and for the backbone rather than relying on entangled mappings.
+Analysis.
+Unlike mixture-of-experts (MoEs) in modern language models (shazeer2017outrageously; xue2024openmoe) designed for efficient sparse parallelism, ours resembles early recommender-system MoEs (ma2018modeling) that capture diverse feature patterns (e.g., rain–rush-hour or fog–low-flow).
+It activates specific experts under local exogenous conditions, resolving inconsistency by decoupling conflicting signals and enabling specialized processing paths (More formal analysis in Appendix A.1).
+Fig.
+4 shows the gating probability distributions across time steps in the AQI-19 experiment, highlighting the smooth composition of expert allocation.
+Empirically, we also validate the module’s efficacy across diverse scenarios in Section.
+5.5.
+4.2 Balance: Dual-Branch Adaptive Fusion
+The balancing stage addresses unbalanced type effects arising from the distributional discrepancies between past and future exogenous contexts.
+To effectively fuse these heterogeneous signals, we employ parallel Siamese networks for independent ST feature extraction, coupled with a context-aware balance mechanism for adaptive fusion.
+Siamese ST Encoders.
+To process the distinct characteristics of past- and future-conditioned semantic representations, we employ a Siamese architecture comprising parallel encoders.
+Each encoder , adaptable to arbitrary spatio-temporal backbones (yu2017spatio; li2017diffusion; guo2019attention; wu2019graph; Bai2020Adaptive; Gao2022On; Satorras2022Multivariate), captures intrinsic spatial correlations and temporal dependencies via:
+where encode high-level spatio-temporal features derived from past and future exogenous contexts, respectively, facilitating subsequent fusion.
+Context-Aware Balancer.
+To address the limitations of static fusion strategies, we employ a dynamic context-aware mechanism to integrate and .
+We first construct a unified context tensor , from which an instance-specific balancing weight is derived via:
+Here, aggregates spatial information to generate a channel descriptor, which passes through an MLP bottleneck with Sigmoid activation .
+The final prediction is synthesized using a residual connection to facilitate information propagation:
+This adaptive calibration effectively prioritizes past versus future contexts based on instance-specific dynamics, yielding a robust and balanced prediction.
+Analysis.
+Prior modeling of past and future exogenous variables typically relies on static fusion, employing either fixed linear weights or global affine transformations optimized over the training distribution (das2023long; giganti2024back; Arango2025ChronosX).
+In contrast, our context-aware mechanism utilizes input-dependent non-linear mappings to approximate arbitrary fusion functions, enabling instance-specific adaptation during inference (More formal analysis in Appendix A.2).
+Fig.
+4 visualizes the weight distributions across inference samples in the AQI-19 experiment; unlike the scattered and irregular patterns observed in baselines, ExoST exhibits smoother and more uniform distributions, underscoring its superior dynamic adaptability.
+Empirically, we also validate the effectiveness of this design in Section.
+5.5.
+5 Experiments
+In this section, we conduct extensive experiments to investigate the following core research questions:
+RQ1: (Universality) Can ExoST achieve consistent improvements on any existing ST backbones?
+RQ2: (Effectiveness) Can ExoST outperform existing other methods across various task scenarios?
+RQ3: (Robustness) Can ExoST maintain its performance when exogenous variables are missing or noisy?
+RQ4: (Mechanism & Rationality) How does ExoST work? Which components or strategies are crucial, and are they sensitive to different designs or hyper-parameters?
+RQ5: (Efficiency & Lightweight) How does ExoST compare in efficiency and parameter count to others?
+5.1 Experimental Setup
 Datasets.
-We conducted various experiments on two public real-world spatio-temporal datasets 111https://zenodo.org/records/7308425, which contains hourly measurements collected from monitoring stations in Madrid during the periods 1 January to 30 June 2019 and 1 January to 30 June 2022.
-Each dataset includes monitoring data from air quality, meteorological, and traffic systems.
-We consider four forecasting tasks: (i) AQI-19 and AQI-22: predict concentrations in air quality system using 2019 / 2022 data, with traffic and meteorological indicators as exogenous variables; (ii) Speed-19 and Intensity-22: predict traffic speed / intensity in the traffic system using 2019 / 2022 data, with other system indicators as exogenous variables.
-All dataset are chronologically split into training, validation, and test sets in a 7:2:1 ratio.
-In addition, it is worth noting that we also include the coding of dates as both past and future exogenous variables as input.
-More details on dataset, please refer to Appendix A.1.
+We evaluate on several public real-world spatio-temporal datasets 111https://zenodo.org/records/7308425, containing hourly measurements from air quality, meteorological, and traffic monitoring stations in Madrid (1 January–30 June, 2019/2022).
+We study four forecasting tasks: (i) AQI-19 and AQI-22: predict in air quality domain concentrations with traffic and meteorological domain variables as exogenous inputs; (ii) Speed-19 and Intensity-22: predict traffic speed / intensity using other system indicators as exogenous variables.
+Datasets are chronologically split into train/val/test with ratio 7:2:1.
+Date encodings are included as past and future exogenous variables.
+More dataset details are provided in Appendix B.1.
 Baseline.
-For performance comparison, we categorize existing baselines into three groups: (i) time series models with exogenous variable modeling : TimeXer (Wang et al., 2024), TiDE (Das et al., 2023), and NBEATSx (Olivares et al., 2023); (ii) spatio-temporal models with exogenous variable integration: MAGCRN (Giganti et al., 2024); (iii) general exogenous variable modeling frameworks: ChronosX (Arango et al., 2025), which allows for the substitution of any backbone.
-Additionally, to evaluate ExoST’s universality, we extensively utilize various classic spatio-temporal backbones, including AGCRN (Bai et al., 2020), GWNet (Wu et al., 2019), GGNN (Satorras et al., 2022), GRUGCN (Gao and Ribeiro, 2022), STGCN (Yu et al., 2017), and DCRNN (Li et al., 2017).
-More details on methods, see Appendix A.2.
+We compared with three comparable types of baselines in the literature: (i) time-series models with exogenous modeling: TiDE (das2023long), NBEATSx (olivares2023neural), TimeXer (wang2024timexer), CrossLinear (zhou2025crosslinear), DAG (qiu2025dagdualcausalnetwork); (ii) spatio-temporal models with exogenous integration: MAGCRN (giganti2024back); (iii) general exogenous modeling framework (for foundation model): ChronosX (Arango2025ChronosX).
+To test ExoST’s universality, we further instantiate it with various classic spatio-temporal backbones, including AGCRN (Bai2020Adaptive), GWNet (wu2019graph), GGNN (Satorras2022Multivariate), GRUGCN (Gao2022On), STGCN (yu2017spatio), and DCRNN (li2017diffusion).
+More details on baseline methods and backbone networks, please refer to Appendix B.2.
 Protocol.
-By default, we use the previous 24 steps of historical data to predict data for the next 24 steps (i.e., 1-day), and evaluate using mean absolute error (MAE), root mean square error (RMSE), mean relative error (MRE), and mean absolute percentage error (MAPE).
-Moreover, to measure the long-term generalization of the model in real-world deployments, we also perform rolling evaluations over 2 and 3 days without retraining.
-For the baselines, we use the parameter settings recommended in their papers and fine-tune them.
-For ExoST, we set the number of gated experts to 4 and the hidden size to 64.
-To ensure fairness, each experiment is repeated five times, and results are reported as the mean standard deviation.
-Best results are highlighted in bold pink and the second-best in underlined blue.
-For more details on the protocol, see Appendix A.3.
-4.2. Effectiveness Study (RQ1)
-Table LABEL:tab:rq1_overall_performance shows the performance comparison of our model with various state-of-the-art methods for time series forecasting or spatio-temporal forecasting that utilize exogenous variable modeling.
-It is worth noting that the first three time series forecasting methods have coupled model designs, while MAGCRN, ChronosX, and our method all use classical AGCRN(Bai et al., 2020) as the backbone network to ensure a fair comparison.
-Our observations are as follows: ❶ Time series forecasting methods generally achieve the worst results, which is consistent with common spatio-temporal forecasting benchmarks (Shao et al., 2024).
-This is because they generally do not consider spatial dependencies, resulting in inaccurate predictions.
-❷ The ChronosX method, which utilizes a spatio-temporal backbone, despite achieving relative progress, often encounters non-convergence issues on some tasks during training (i.e., Speed-19 and Intensity-22).
-Despite repeated parameter tuning, this issue cannot be avoided.
-We attribute this to the design of the underlying Chronos-type (Ansari et al., [n.
-d.]) time series foundation model.
-❸ MAGCRN often achieves suboptimal results because it is specifically designed for spatio-temporal prediction scenarios and still ignores the inconsistent and unbalanced effects of exogenous variables.
-❹ Our ExoST achieves competitive performance across all metrics on the four tasks.
-By effectively selecting and balancing exogenous information, the ExoST can adapt to a variety of input sources, maintaining stable performance.
-4.3. Universality Study (RQ2)
-We further aim to demonstrate the universality of our ExoST’s “select, then balance” paradigm across various spatio-temporal backbones.
-Table LABEL:tab:rq1_3day illustrates the performance improvements of our proposed ExoST for 3-day predictions across four tasks and six models.
-The ✗ column reflects the the performance metrics of the backbone model alone on the target system, while the ✓ column displays the improved results achieved by our ExoST through exogenous variable modeling.
-The relative improvement is quantified in the column.
-Based on these results, we make the following observations: ❶ Our ExoST exhibits significant performance improvements across different architectures and datasets, attributed to an effective exogenous variable modeling paradigm.
-❷ From the model perspective, the extent of performance improvement varies among different backbone networks, with AGCRN yielding the optimal overall enhancement; however, even the least performing Gwnet model achieved substantial performance gains.
-❸ From the task perspective, air quality prediction consistently showed stable performance improvements, as exogenous information like traffic and meteorology significantly impacts it; conversely, traffic speed prediction saw only moderate improvements due to the limited influence of exogenous information such as air quality.
-We further supplement results for 1- and 2-day predictions in Appendix B.1 and additionally find that ❹ as the prediction horizon extends, our ExoST achieves more significant gains, which also indicates the strong generalization capability of our method.
-4.4. Robustness Study (RQ3)
-In complex real-world scenarios, exogenous variable data in spatio-temporal systems may not always be collected reliably (for example, signal loss caused by sensor anomalies or maintenance), which can affect prediction accuracy and model robustness.
-To evaluate the robustness of the ExoST under such adverse conditions, we designed a systematic experiment to simulate real-world data corruption scenarios.
-Specifically, we randomly replaced exogenous variables with varying degrees of degradation (20%, 40%, 60%, and 80%) with zero values or randomly generated values from a standard normal distribution to evaluate the model’s ability to maintain predictive performance despite exogenous information corruption.
-The 3-day prediction experiment results shown in Table LABEL:tab:mask_3day reveal several interesting phenomena: ❶ As the proportion of missing data increases, model performance gradually decreases in most cases.
-However, the overall absolute performance drop is not significant, indicating that our framework can still maintain reasonable performance with limited auxiliary information from exogenous variables.
-❷ For some tasks, appropriate missing signals can lead to improved performance, primarily because random data corruption can be viewed as a simple data augmentation strategy.
-By introducing different data perturbations, the model is forced to learn to predict under a variety of incomplete or noisy data conditions, thereby enhancing its generalization and robustness.
-We also provide a performance comparison for the 1-day and 2-day scenarios in Appendix B.2, observing similar results.
-4.5. Mechanism Study (RQ4)
-4.5.1. Ablation Study
-We first conducted an ablation study to systematically evaluate the impact of different components on ExoST from two perspectives: data input and model component.
-From the data perspective, we aimed to investigate the auxiliary role of different types of exogenous variables in the final prediction.
-Specifically, due to the universality of date information, here, we treated it as a separate category.
-As shown in Table LABEL:tab:performance_1day, we observed that ❶ future exogenous variables outperformed past, indicating that forward-looking information provides more direct predictive signals.
-However, combining both past and future yielded better performance than using either type alone, which suggests that the combination can provide complementary patterns.
-❷ We also found that date exogenous variables performed poorly when used in isolation but led to improvements when combined with any other type.
-❸ Best performance was achieved by combining all exogenous variables.
-This demonstrates that ExoST can effectively exploit different information, extracting and balancing contributions from different temporal perspectives and contextual information.
-From the model perspective, we aimed to investigate the auxiliary role of different module components in the final prediction.
-We constructed two variants by removing the MoE Selector and the Context-aware Balancer (achieved through direct element-wise fusion), respectively.
-As Figure 9 shows, ❶ w/ Selector variant significantly degrades performance in all metrics, demonstrating the critical importance of addressing inconsistent variable effects.
-Performance degradation indicates that without proper semantic alignment and adaptive selection mechanisms, the model struggles to disentangle heterogeneous exogenous variables from endogenous variables, leading to feature space competition and channel entanglement issues.
-❷ w/ Balancer variant causes even more pronounced performance degradation than w/ Selector variant, highlighting that temporal distribution differences between past and future exogenous information pose a more fundamental challenge, requiring sophisticated balancing mechanisms to handle noise-contaminated past information and uncertainty-amplified future forecasts.
-More experimental results are shown in Appendix B.3.1.
-4.5.2. Parameter Study.
-We then conducted a parameter study to evaluate the impact of different parameter on ExoST.
-As Figure 5 shows, we can observe that ❶ The hidden size of 64 achieves optimal performance in all prediction horizons, with larger sizes leading to performance degradation due to overfitting.
-❷ Expert number of 4 provides the best balance between model capacity and computational efficiency.
-Too few experts (one or two) fail to capture the diverse semantic patterns needed for disentangling inconsistent variable effects, while excessive experts (eight) lead to overselection where redundant experts compete for similar patterns, diminishing the effectiveness of the mixture-of-experts gating mechanism designed to mitigate channel entanglement.
-These observations validate the effectiveness of our lightweight design choices in the latent-space gated expert module.
-4.5.3. Strategy Study.
-We also use several alternative strategies to test the rationality of our design.
-These include: (i) Shared Parameter: Two siamese branches share parameters to learn a unified spatio-temporal representation.
-(ii) Simple Weight: Adopts a predefined balancing weight () to fuse the outputs from two siamese branches.
-(iii) Learnable Weight: Dynamically adjusts the relative importance of past and future information through learnable parameters.
-(iv) Attention: Utilizes cross-attention mechanisms to dynamically balance the contributions from two siamese branches.
-As shown in Figure 6, we observe that: ❶ ExoST achieves competitive performance in other strategies, demonstrating the effectiveness of our design choice.
-❷ Simple Weight shows significant performance degradation due to its static balancing strategy that does not adapt to varying input contexts.
-❸ Attention and Learnable Weight strategies underperform while increasing computational complexity, as attention mechanisms introduce excessive parameter overhead and learnable weights lack the sophisticated context modeling capabilities needed to handle nuanced differences between exogenous variables of the past and future.
-More experimental results are shown in Appendix B.3.2.
-4.6. Efficiency & Lightweight Study (RQ5)
-We compared the differences in training time, memory footprint, and performance between ExoST and other time series or spatio-temporal forecasting models that incorporate exogenous variables on the AQI-19 dataset.
-Due to varying convergence epochs, for a fair comparison, we only present the average training time for the first 100 epochs.
-As shown in Figure 7, compared to several other methods, ExoST demonstrates a balance of performance, efficiency, and model parameter count.
-Furthermore, we can observe that: ❶ compared to time series models with different backbones, despite its lightweight model design, it leads to poor performance.
-❷ compared to spatio-temporal method and the general exogenous variables modeling method with the same backbone (i.e., MAGCRN, ChronosX), we achieved further performance improvement while maintaining similar parameters and training time.
-5. Conclusion and Future Work
-This paper investigates the challenges of modeling exogenous variables for spatio-temporal forecasting and explores effective implementation options.
-We propose a novel framework, ExoST, which follows a “select, then balance” paradigm.
-It first dynamically selects and restructures salient signals via a latent space gated expert module to address inconsistent variable effects.
-It then incorporates a dual-branch architecture with a context-aware weighting mechanism to address imbalanced type effects.
+We use the past 24 steps to forecast the next 24 steps (i.e., 1-day) and report MAE, RMSE, MAPE, and MRE.
+To assess long-term robustness, we additionally perform 2- and 3-day rolling forecasts without retraining.
+Baseline models follow their recommended settings.
+For ExoST, we use 4 gated experts and hidden size 64.
+Each experiment is run five times; we report mean standard deviation.
+Best results appear in bold blue, second-best in underlined green.
+Full protocol details can be found in Appendix B.3.
+5.2 Universality Study (RQ1)
+We first evaluate the universality of the “select, then balance” paradigm in ExoST across diverse ST backbones.
+Figure 5 illustrates the absolute and average relative performance improvements of ExoST across four tasks (left) and six models (right) for 3-day forecasting.
+❶ ExoST yields consistent performance boosts across varying architectures and datasets, validating the efficacy of our exogenous modeling paradigm.
+❷ From the task perspective, air quality forecasting achieves substantial gains due to the strong influence of exogenous factors (e.g., traffic and meteorology), whereas traffic speed forecasting sees moderate improvements, reflecting a limited dependency on exogenous variables.
+❸ From the model perspective, while gains vary across backbones, most exceed 20%; notably, even the vanilla GWNet achieves a significant improvement of nearly 10%.
+❹ Supplementary results for 1-day and 2-day horizons (Tables C.2 and C.3 in Appendix C.1) further reveal that performance gains amplify as the forecasting horizon extends, underscoring the robust generalization capability of our approach.
+| Datasets | AQI-19 | Speed-19 ∗ | AQI-22 | Intensity-22 | |||||||||
+| Method | Metric | 1-day | 2-day | 3-day | 1-day | 2-day | 3-day | 1-day | 2-day | 3-day | 1-day | 2-day | 3-day |
+| TiDE (das2023long) | MAE | ||||||||||||
+| RMSE | |||||||||||||
+| MAPE (%) | |||||||||||||
+| MRE (%) | |||||||||||||
+| NBEATSx (olivares2023neural) | MAE | ||||||||||||
+| RMSE | |||||||||||||
+| MAPE (%) | |||||||||||||
+| MRE (%) | |||||||||||||
+| TimeXer (wang2024timexer) | MAE | ||||||||||||
+| RMSE | |||||||||||||
+| MAPE (%) | |||||||||||||
+| MRE (%) | |||||||||||||
+| CrossLinear (zhou2025crosslinear) | MAE | ||||||||||||
+| RMSE | |||||||||||||
+| MAPE (%) | |||||||||||||
+| MRE (%) | |||||||||||||
+| DAG (qiu2025dagdualcausalnetwork) | MAE | ||||||||||||
+| RMSE | |||||||||||||
+| MAPE (%) | |||||||||||||
+| MRE (%) | |||||||||||||
+| MAGCRN (giganti2024back) | MAE | ||||||||||||
+| RMSE | 108.66±17.87 | ||||||||||||
+| MAPE (%) | 42.53±0.41 | 43.79±4.21 | 45.89±4.35 | 49.74±4.19 | |||||||||
+| MRE (%) | |||||||||||||
+| ChronosX (Arango2025ChronosX) | MAE | ||||||||||||
+| RMSE | |||||||||||||
+| MAPE (%) | |||||||||||||
+| MRE (%) | |||||||||||||
+| ExoST | MAE | 9.33±0.13 | 10.04±0.19 | 10.41±0.12 | 3.79±0.01 | 3.80±0.01 | 3.80±0.01 | 7.33±0.21 | 7.61±0.23 | 7.94±0.29 | 31.91±1.35 | 50.59±0.04 | 64.13±3.86 |
+| RMSE | 14.24±0.18 | 15.12±0.26 | 15.51±0.16 | 31.73±0.02 | 31.78±0.02 | 31.85±0.02 | 11.14±0.21 | 11.52±0.49 | 11.79±0.51 | 56.49±2.54 | 89.10±7.36 | ||
+| MAPE (%) | 38.57±0.32 | 44.65±0.08 | 9.30±0.03 | 9.32±0.03 | 9.31±0.03 | 17.63±0.97 | 26.68±1.69 | 34.52±2.90 | |||||
+| MRE (%) | 36.09±0.49 | 38.84±0.75 | 40.34±0.47 | 9.88±0.03 | 9.89±0.03 | 9.90±0.03 | 39.23±1.11 | 40.74±1.24 | 42.50±1.57 | 11.57±0.49 | 18.31±1.14 | 23.16±1.40 | |
+5.3 Effectiveness Study (RQ2)
+Table 2 presents a performance comparison of our ExoST against state-of-the-art time-series, spatio-temporal, and foundation models with exogenous variable modeling techniques.
+To ensure a fair comparison, we employ the standard AGCRN (Bai2020Adaptive) as the backbone for all methods utilizing a decoupled design.
+Our observations are as follows: ❶ Pure time-series methods (first 5) consistently yield the lowest performance, aligning with established ST benchmarks (shao2024exploring).
+This deficiency stems from their inability to model spatial dependencies, resulting in suboptimal accuracy.
+❷ While ChronosX achieves relative gains by integrating the spatio-temporal backbone, it frequently suffers from non-convergence on specific tasks (e.g., Speed-19 and Intensity-22), despite extensive hyperparameter tuning.
+We attribute this instability to the inherent design of the underlying Chronos-type foundation models (ansarichronos).
+❸ MAGCRN often produces suboptimal results as it is tailored for specific scenarios yet neglects the inconsistent and unbalanced effects of exogenous variables.
+❹ Conversely, our ExoST achieves superior performance across all metrics on the four tasks.
+By effectively selecting and balancing exogenous information, ExoST adaptively handles diverse input sources to maintain robust performance.
+5.4 Robustness Study (RQ3)
+In real-world scenarios, spatio-temporal systems frequently encounter unreliable exogenous data collection (e.g., signal loss from sensor anomalies), which compromises prediction accuracy and robustness.
+To assess the ExoST under such conditions, we designed a systematic experiment simulating data corruption.
+Specifically, we randomly replaced exogenous variables at varying degradation levels (20%, 40%, 60%, and 80%) with zeros or noise from a standard normal distribution .
+The 3-day prediction results in Table 3 reveal that: ❶ While performance gradually declines as the missing data ratio increases, the absolute drop remains minimal, indicating our framework maintains reasonable efficacy even with limited auxiliary information.
+❷ Surprisingly, for certain tasks, appropriate levels of missing signals improve performance.
+This suggests that random corruption acts as a data augmentation strategy, forcing the model to predict under incomplete or noisy conditions, thereby enhancing generalization and robustness.
+Similar trends for 1-day and 2-day scenarios are reported in Appendix C.2.
+| Datasets | AQI-19 | Speed-19∗ | Intensity-22 | ||||||||||
+| Masking Strategy | MAE | RMSE | MAPE(%) | MRE(%) | MAE | RMSE | MAPE(%) | MRE(%) | MAE | RMSE | MAPE(%) | MRE(%) | |
+| Zero | 20% | 108.23±18.00 | 34.19±3.45 | ||||||||||
+| 40% | |||||||||||||
+| 60% | |||||||||||||
+| 80% | |||||||||||||
+| Random | 20% | ||||||||||||
+| 40% | |||||||||||||
+| 60% | |||||||||||||
+| 80% | |||||||||||||
+| No Masking | 10.41±0.12 | 15.51±0.16 | 42.62±0.52 | 38.84±0.75 | 3.80±0.01 | 31.85±0.02 | 9.31±0.03 | 9.90±0.03 | 64.13±3.86 | 23.16±1.40 | |||
+5.5 Mechanism & Rationality Study (RQ4)
+Ablation Study. We conducted an ablation study to systematically evaluate the impact of different components on ExoST from both data and model perspectives.
+| P | F | D | MAE | RMSE | MAPE (%) | MRE (%) |
+| ✓ | ✓ | ✓ | 9.33±0.13 | 14.24±0.18 | 38.57±0.32 | 36.09±0.49 |
+From the data perspective, we investigated the auxiliary role of distinct exogenous variable types.
+Given the universality of date information, we treat it as a separate category.
+As shown in Table 6: ❶ Future exogenous variables outperform past variables, indicating that forward-looking information provides more direct predictive signals; however, combining both yields superior performance, suggesting they offer complementary patterns.
+❷ Date exogenous variables perform poorly in isolation but drive improvements when combined with other types.
+❸ The best performance is achieved by integrating all exogenous variables, demonstrating that ExoST effectively extracts and balances contributions from diverse temporal contexts.
+From the model perspective, we examined the contribution of specific modules by constructing variants w/o the Gated Expert Selector or the Context-aware Balancer (replaced by direct element-wise fusion).
+As Figure 6 illustrates: ❶ The w/o Selector variant suffers significant degradation across all metrics, proving the importance of addressing inconsistent variable effects.
+Without semantic alignment and adaptive selection, the model struggles to disentangle heterogeneous exogenous variables from endogenous ones, leading to feature space competition.
+❷ The w/o Balancer variant exhibits even more pronounced degradation, highlighting that temporal distribution shifts between past and future exogenous information pose a fundamental challenge.
+Sophisticated balancing mechanisms are key to manage noise-contaminated data and uncertainty-laden future forecasts.
+More experimental results are shown in Appendix C.3.1.
+Parameter Study.
+We then performed a parameter sensitivity analysis to assess the impact of hyper-parameters on ExoST.
+As shown in Figure 8: ❶ A hidden size of 64 achieves optimal performance across all prediction horizons, with larger sizes causing degradation due to overfitting.
+❷ Setting the number of experts to 4 provides the best trade-off between model capacity and computational efficiency.
+Fewer experts fail to capture the diverse semantic patterns required to disentangle inconsistent effects, while excessive experts lead to redundancy and overselection, diminishing the effectiveness of the mixture-of-experts gating mechanism.
+These findings validate the lightweight design choices of our latent-space gated expert module.
+Strategy Study.
+We compared our approach against several alternative fusion strategies to validate our design rationale:(i) Shared Parameter: Siamese branches sharing parameters to learn a unified representation.(ii) Simple Weight: A static balancing weight () fusing outputs.(iii) Learnable Weight: Dynamically adjusting importance via learnable parameters.(iv) Attention: Using cross-attention to balance contributions.As shown in Figure 8: ❶ ExoST outperforms all alternative strategies, confirming the effectiveness of our design.❷ Simple Weight degrades performance significantly due to its inability to adapt to varying input contexts.❸ Attention and Learnable Weight strategies underperform while increasing computational cost; attention mechanisms introduce excessive parameter overhead, and learnable weights lack the sophisticated context modeling required to handle the nuanced differences between past and future exogenous variables.
+More experimental results are shown in Appendix C.3.2.
+5.6 Efficiency & Lightweight Study (RQ5)
+We evaluated the training time, memory footprint, and performance of ExoST against other forecasting models incorporating exogenous variables on the AQI-19 dataset.
+To ensure fair comparison given varying convergence rates, we report the average training time (first 100 epochs).
+As shown in Figure 9, ExoST strikes a favorable balance between performance, efficiency, and parameter count.
+Specifically: ❶ Compared to time series models, purely lightweight designs often lead to inferior performance.
+❷ Compared to ST and general exogenous variable modeling methods using the same backbone (i.e., MAGCRN, ChronosX), our approach achieves superior performance while maintaining comparable parameter counts and training times.
+6 Conclusion and Future Work
+This paper presents the first systematic study of the challenges inherent in modeling exogenous variables for spatio-temporal forecasting and explores effective implementation strategies.
+We propose a simple yet effective general framework, ExoST.
+This framework follows a “select, then balance” paradigm, which effectively addresses the problems of inconsistent variable effects and unbalanced type effects.
 Extensive experiments demonstrate its effectiveness, universality, and robustness.
-In future work, we aim to explore how to broaden the modal availability of exogenous variables and their feasibility in data-scarce scenarios.
-Supplementary material
-Select, then Balance: A Plug-and-Play Framework for Exogenous-Aware Spatio-Temporal Forecasting
-[…]
-[sections] Table of Contents
-[sections]l1
-Appendix A Experimental Details
-A.1. Datasets Details
-The statistics of the two real-world datasets used in this paper are summarized in Table 1.
-| Attribute | Madrid-19 | Madrid-22 | ||||
-| Time range | 01/01/2019 – 30/06/2019 | 01/01/2022 – 30/06/2022 | ||||
-| Frames | 4344 | 4344 | ||||
-| Sampling Rate | 1 hour | 1 hour | ||||
-| Data points | 2.03 M | 2.03 M | ||||
-| Variables | 20 | 20 | ||||
-| Variable Domain |
+In future work, we aim to explore how to expand the modality availability of exogenous variables and their feasibility in extremely data-scarce scenarios.
+Acknowledgment
+This work is mainly supported by the National Natural Science Foundation of China (No.
+62402414).
+This work is also supported by the Guangzhou-HKUST(GZ) Joint Funding Program (No.
+2024A03J0620), Guangzhou Municipal Science and Technology Project (No.
+2023A03J0011), the Guangzhou Industrial Information and Intelligent Key Laboratory Project (No.
+2024A03J0628), and a grant from State Key Laboratory of Resources and Environmental Information System, and Guangdong Provincial Key Lab of Integrated Communication, Sensing and Computation for Ubiquitous Internet of Things (No.
+2023B1212010007).
+Appendix
+Appendix A More Analysis
+A.1 More Analysis on Select Module
+We provide theoretical insights into why the proposed select stage (conditional embedding + latent-space gated experts) improves exogenous-aware spatio-temporal forecasting, addressing the challenge of inconsistent variable effects.
+Our analysis is backbone-agnostic and focuses on representation, approximation, and optimization perspectives.
+✦Input-dependent linear operator and regime-adaptive processing. The gated expert selector in Eq. (4) can be reformulated as an input-dependent linear operator:
+where is a simplex-valued gate.
+Unlike a static projection that enforces a global transformation, Eq.
+(A.1) implements a dynamic filter whose parameters evolve with local spatio-temporal contexts encoded in .
+This explicitly models the inconsistent effects, where the contribution of endogenous channels varies across different weather conditions or temporal regimes.
+✦Latent regime approximation.
+We postulate that the data-generating process involves a continuous latent regime variable (e.g., a continuous spectrum from clear to extreme weather).
+The predictive distribution can be viewed as an integral over these regimes:
+Our module approximates this integral via discretization.
+The conditional embedding aligns heterogeneous sources, enabling the gate to approximate the regime posterior .
+The experts then serve as support points in the function space, effectively decomposing the complex multimodal mapping into a mixture of local experts:
+✦Reduced gradient interference via soft routing.
+Inconsistent effects also manifest as gradient conflicts: samples from different contexts induce updates in opposing directions when sharing parameters.
+For an expert parameter (e.g., ), the per-sample gradient satisfies
+where is the gate weight for sample .
+Thus, samples primarily update the experts they activate, which decomposes conflicting gradients across experts and reduces negative transfer.
+Empirically, this is consistent with smoother gate trajectories and clearer expert specialization (Fig.
+4), improving optimization stability and downstream forecasting accuracy.
+Overall, the key advantage of the Select Module is that it explicitly incorporates exogenous-condition-induced variations in variable effects into both the representation learning and routing mechanisms.
+Specifically, the conditional embedding first fuses endogenous histories with heterogeneous, multi-source exogenous signals into context-dependent latent representations, thereby establishing a unified semantic basis for subsequent selection.
+Building on this latent space, the gated experts further instantiate an input-dependent dynamic operator, enabling the model to automatically activate more appropriate processing pathways across different time steps, spatial nodes, and exogenous contexts.
+Meanwhile, soft routing assigns the learning signal from each contextual sample primarily to its corresponding expert parameters, effectively alleviating gradient interference and negative transfer that commonly arise under shared-parameter training, which in turn promotes context-specialized expert division of labor and improves optimization efficiency.
+Consequently, by combining conditional representation with expert selection, the Select stage realizes adaptive modeling of inconsistent variable effects and provides the downstream backbone with more robust and controllable input features.
+A.2 More Analysis on Balance Module
+Although existing methods (wang2024timexer; giganti2024back; Arango2025ChronosX) have explored various strategies for balancing past and future exogenous variables, our approach introduces a fundamental distinction in its design philosophy.
+We theoretically scrutinize the discrepancies among these methods through the lens of functional complexity and the representational capacity of their respective function spaces:
+Fixed Weighting Strategy: This approach operates as a rudimentary static linear mapping, formalized as:
+The representational capacity of this method is severely constrained, as the dimension of the function space is , governed exclusively by the scalar hyperparameter .
+It imposes a rigid interpolation that is agnostic to data distribution.
+Learnable Weighting Strategy: While this method introduces optimization, it fundamentally remains a linear mapping within the family of affine transformations:
+Although the parameters are learnable, the objective is to converge upon a single, global optimal weighted balance across the entire training set.
+Consequently, the learned weights remain static during the inference phase and cannot adapt to the specific variations of individual samples.
+Context-Aware Balancing (Ours): In stark contrast, our proposed mechanism functions as an input-dependent, non-linear mapping:
+Theoretically, this formulation is capable of approximating arbitrary fusion functions.
+By conditioning the weights on the input itself, the model achieves context sensitivity at inference time, allowing for the dynamic modulation of the fusion strategy based on immediate data characteristics.
+Therefore, our methodology diverges significantly from prior static strategies by offering superior dynamic adaptability.
+To validate this theoretical advantage empirically, Section 5.5 presents a comprehensive ablation study comparing our approach against degenerate fixed-weight baselines, learnable-weight strategies, and attention-based mechanisms.
+The empirical results robustly demonstrate the efficacy and necessity of our context-aware balancing strategy.
+Appendix B Experimental Details
+B.1 Datasets Details
+The statistics of the two real-world datasets used in this paper are summarized in Table B.1.
+| Attribute | Madrid-19 | Madrid-22 |
+| Time range | 01/01/2019 – 30/06/2019 | 01/01/2022 – 30/06/2022 |
+| Frames | 4344 | 4344 |
+| Sampling Rate | 1 hour | 1 hour |
+| Data points | 2.03 M | 2.03 M |
+| Variables | 20 | 20 |
+| Variable Domain | Air quality, Traffic, Meteorological, Date | Air quality, Traffic, Meteorological, Date |
 Graph Construction for GNN-based Models.
@@
 These temporal features are added as universal exogenous variables to four tasks.
-A.2. Baseline Details
+B.2 Baseline Details
 In this appendix, we provide detailed descriptions of the advanced exogenous variable modeling methods and different spatio-temporal backbone models used in our default evaluation.
-A.2.1. Exogenous Variable Modeling Methods
+B.2.1 Exogenous Variable Modeling Methods
 These methods are grouped into three main categories based on their approach to handling exogenous variables.
 TS Models w/ Exogenous Variable Integration.
-TiDE (Das et al., 2023): TiDE is an encoder–decoder model built on multilayer perceptrons (MLPs) that fuses historical time-series data and covariates via dense MLP residual blocks to map them into low-dimensional representations, then generates forecast sequences by incorporating future covariates.
+TiDE (das2023long): TiDE is an encoder–decoder model built on multilayer perceptrons (MLPs) that fuses historical time-series data and covariates via dense MLP residual blocks to map them into low-dimensional representations, then generates forecast sequences by incorporating future covariates.
 https://github.com/google-research/google-research/tree/master/tide
-TimeXer (Wang et al., 2024): TimeXer is a Transformer model tailored for time series forecasting with exogenous variables: it captures temporal dependencies in the endogenous series via patch-level self-attention and effectively integrates exogenous information through variable-level cross attention and learnable global endogenous tokens to produce high-quality predictions of the target series.
+TimeXer (wang2024timexer): TimeXer is a Transformer model tailored for time series forecasting with exogenous variables: it captures temporal dependencies in the endogenous series via patch-level self-attention and effectively integrates exogenous information through variable-level cross attention and learnable global endogenous tokens to produce high-quality predictions of the target series.
 https://github.com/thuml/TimeXer
-NBEATSx (Olivares et al., 2023): NBEATSx extends the original N-BEATS by incorporating convolutional encoding substructures for time-varying and static exogenous variables within a hierarchical stacked residual block architecture, enabling interpretable decomposition of trend, seasonality, and exogenous factors for high-quality target-series forecasting.
+NBEATSx (olivares2023neural): NBEATSx extends the original N-BEATS by incorporating convolutional encoding substructures for time-varying and static exogenous variables within a hierarchical stacked residual block architecture, enabling interpretable decomposition of trend, seasonality, and exogenous factors for high-quality target-series forecasting.
 https://github.com/cchallu/nbeatsx
+CrossLinear (zhou2025crosslinear):CrossLinear is a Linear-based model that introduces a plug-and-play cross-correlation embedding module to capture time-invariant dependencies between endogenous and exogenous variables, utilizing patch-wise processing and a global linear head to effectively model both short-term and long-term temporal patterns for precise forecasting.
+https://github.com/mumiao2000/CrossLinear
+DAG (qiu2025dagdualcausalnetwork): DAG is a general framework that utilizes a Dual Causal Network along temporal and channel dimensions, it explicitly models how historical exogenous variables affect future exogenous variables (temporal) and historical endogenous variables (channel), and then injects these discovered causal relationships to improve the prediction of future endogenous variables.
+https://github.com/decisionintelligence/DAG
 ST Models w/ Exogenous Variable Integration.
-MAGCRN (Giganti et al., 2024): MAGCRN is a spatio-temporal model that extends AGCRN into a dual-path architecture, using separate AGCRN blocks to encode historical observations and future exogenous covariates.
+MAGCRN (giganti2024back): MAGCRN is a spatio-temporal model that extends AGCRN into a dual-path architecture, using separate AGCRN blocks to encode historical observations and future exogenous covariates.
 https://github.com/polimi-ispl/MAGCRN
 General Exogenous Variable Integration Framework.
-ChronosX (Arango et al., 2025): ChronosX extends MAGCRN by integrating two exogenous adaptation modules: Instance-wise Interpolation Block (IIB) and Offset Integration Block (OIB).
+ChronosX (Arango2025ChronosX): ChronosX extends MAGCRN by integrating two exogenous adaptation modules: Instance-wise Interpolation Block (IIB) and Offset Integration Block (OIB).
 IIB adaptively interpolates exogenous inputs across time steps, while OIB injects time-dependent bias offsets into the output layer, enabling more flexible integration of future covariates.
@@
 https://github.com/amazon-science/chronos-forecasting/tree/chronosx
-A.2.2. Spatio-Temporal Backbone Models
+B.2.2 Spatio-Temporal Backbone Models
 To ensure fair use of these models, we use a popular spatio-temporal forecasting benchmark library: https://github.com/TorchSpatiotemporal/tsl.
-AGCRN (Bai et al., 2020): AGCRN is an adaptive graph-based recurrent model that captures fine-grained spatio-temporal dependencies by jointly learning node-specific patterns and data-driven graph structures, eliminating the need for pre-defined adjacency matrices.
-GWNet (Wu et al., 2019): GWNet is a spatial-temporal model that combines diffusion-based graph convolutions with dilated causal temporal convolutions, and incorporates a learnable self-adaptive adjacency matrix to capture latent spatial relationships from data.
-GGNN (Satorras et al., 2022): GGNN is a recurrent graph-based model that applies gated update mechanisms to iteratively refine node representations through message passing, enabling effective modeling of long-range structural dependencies.
-GRUGCN (Gao and Ribeiro, 2022): GRUGCN is a spatio-temporal neural architecture that integrates graph convolutions for spatial feature extraction and gated recurrent units for sequential modeling, effectively capturing dynamic dependencies in traffic forecasting scenarios.
-STGCN (Yu et al., 2017): STGCN is a graph-based temporal forecasting model that integrates graph convolutional layers with temporal convolution modules to jointly learn spatial and temporal dependencies in structured time series data.
-DCRNN (Li et al., 2017): DCRNN is a graph-based recurrent forecasting model that captures spatial dependencies via diffusion convolutions on directed graphs and models temporal dynamics through a sequence-to-sequence recurrent architecture.
-A.3. Protocol Details
-A.3.1. Metrics Detail.
+AGCRN (Bai2020Adaptive): AGCRN is an adaptive graph-based recurrent model that captures fine-grained spatio-temporal dependencies by jointly learning node-specific patterns and data-driven graph structures, eliminating the need for pre-defined adjacency matrices.
+GWNet (wu2019graph): GWNet is a spatial-temporal model that combines diffusion-based graph convolutions with dilated causal temporal convolutions, and incorporates a learnable self-adaptive adjacency matrix to capture latent spatial relationships from data.
+GGNN (Satorras2022Multivariate): GGNN is a recurrent graph-based model that applies gated update mechanisms to iteratively refine node representations through message passing, enabling effective modeling of long-range structural dependencies.
+GRUGCN (Gao2022On): GRUGCN is a spatio-temporal neural architecture that integrates graph convolutions for spatial feature extraction and gated recurrent units for sequential modeling, effectively capturing dynamic dependencies in traffic forecasting scenarios.
+STGCN (yu2017spatio): STGCN is a graph-based temporal forecasting model that integrates graph convolutional layers with temporal convolution modules to jointly learn spatial and temporal dependencies in structured time series data.
+DCRNN (li2017diffusion): DCRNN is a graph-based recurrent forecasting model that captures spatial dependencies via diffusion convolutions on directed graphs and models temporal dynamics through a sequence-to-sequence recurrent architecture.
+B.3 Protocol Details
+B.3.1 Metrics Detail.
 We use different metrics such as MAE, RMSE, MRE, and MAPE. Formally, these metrics are formulated as follows:
@@
 The mean relative error is defined as the MAE divided by the norm of the target window.
-A.3.2. Parameter Detail.
+B.3.2 Parameter Detail.
 The model is trained for 500 epochs using the AdamW optimizer with a cosine annealing learning rate scheduler.
@@
 All experiments are carried out on a Linux server equipped with a 1 AMD EPYC 7763 128-Core Processor CPU with 256GB memory and 4 NVIDIA RTX A6000 GPUs, each with 48GB memory.
-Appendix B More Results
-B.1. More Result on Universality Study
+Appendix C More Results
+C.1 More Result on Universality Study
 | Model | AGCRN | GWNet | GGNN | GRUGCN | STGCN | DCRNN | |||||||||||||
 | w/ ExoST | ✗ | ✓ | (%) | ✗ | ✓ | (%) | ✗ | ✓ | (%) | ✗ | ✓ | (%) | ✗ | ✓ | (%) | ✗ | ✓ | (%) | |
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+| AQI-19 | MAE | 10.41±0.12 | |||||||||||||||||
+| RMSE | 15.51±0.16 | ||||||||||||||||||
+| MAPE(%) | 44.65±0.08 | ||||||||||||||||||
+| MRE(%) | 40.34±0.47 | ||||||||||||||||||
+| Speed-19 | MAE | 3.80±0.01 | |||||||||||||||||
+| RMSE | 31.85±0.02 | ||||||||||||||||||
+| MAPE(%) | 9.31±0.03 | ||||||||||||||||||
+| MRE(%) | 9.90±0.03 | ||||||||||||||||||
+| AQI-22 | MAE | 7.94±0.29 | |||||||||||||||||
+| RMSE | 11.79±0.51 | ||||||||||||||||||
+| MAPE(%) | 50.27±1.18 | ||||||||||||||||||
+| MRE(%) | 41.18±1.83 | ||||||||||||||||||
+| Intensity-22 | MAE | 61.87±3.80 | |||||||||||||||||
+| RMSE | 105.44±5.53 | ||||||||||||||||||
+| MAPE(%) | 31.74±1.69 | ||||||||||||||||||
+| MRE(%) | 22.35±1.37 | ||||||||||||||||||
 | Model | AGCRN | GWNet | GGNN | GRUGCN | STGCN | DCRNN | |||||||||||||
 | w/ ExoST | ✗ | ✓ | (%) | ✗ | ✓ | (%) | ✗ | ✓ | (%) | ✗ | ✓ | (%) | ✗ | ✓ | (%) | ✗ | ✓ | (%) | |
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-We provide more experimental results on universality (i.e., for 1-day and 2-day scenarios), as shown in Tables 2 and 3.
+| AQI-19 | MAE | 10.04±0.19 | |||||||||||||||||
+| RMSE | 15.12±0.26 | ||||||||||||||||||
+| MAPE(%) | 42.62±0.52 | ||||||||||||||||||
+| MRE(%) | 38.84±0.75 | ||||||||||||||||||
+| Speed-19 | MAE | 3.77±0.06 | |||||||||||||||||
+| RMSE | 31.38±0.02 | ||||||||||||||||||
+| MAPE(%) | 9.05±0.23 | ||||||||||||||||||
+| MRE(%) | 9.83±0.15 | ||||||||||||||||||
+| AQI-22 | MAE | 7.35±0.21 | |||||||||||||||||
+| RMSE | 11.47±0.39 | ||||||||||||||||||
+| MAPE(%) | 46.59±0.64 | ||||||||||||||||||
+| MRE(%) | 39.35±1.11 | ||||||||||||||||||
+| Intensity-22 | MAE | 50.53±3.94 | |||||||||||||||||
+| RMSE | 89.10±7.36 | ||||||||||||||||||
+| MAPE(%) | 26.11±2.67 | ||||||||||||||||||
+| MRE(%) | 18.29±1.43 | ||||||||||||||||||
+| Model | AGCRN | GWNet | GGNN | GRUGCN | STGCN | DCRNN | |||||||||||||
+| w/ ExoST | ✗ | ✓ | (%) | ✗ | ✓ | (%) | ✗ | ✓ | (%) | ✗ | ✓ | (%) | ✗ | ✓ | (%) | ✗ | ✓ | (%) | |
+| AQI-19 | MAE | 9.33±0.13 | |||||||||||||||||
+| RMSE | 14.24±0.18 | ||||||||||||||||||
+| MAPE(%) | 38.57±0.32 | ||||||||||||||||||
+| MRE(%) | 35.85±0.71 | ||||||||||||||||||
+| Speed-19 | MAE | 3.70±0.06 | |||||||||||||||||
+| RMSE | 30.92±0.04 | ||||||||||||||||||
+| MAPE(%) | 8.59±0.21 | ||||||||||||||||||
+| MRE(%) | 9.64±0.16 | ||||||||||||||||||
+| AQI-22 | MAE | 6.81±0.30 | |||||||||||||||||
+| RMSE | 10.72±0.58 | ||||||||||||||||||
+| MAPE(%) | 31.80±0.70 | ||||||||||||||||||
+| MRE(%) | 36.44±1.62 | ||||||||||||||||||
+| Intensity-22 | MAE | 29.74±1.21 | |||||||||||||||||
+| RMSE | 48.88±1.95 | ||||||||||||||||||
+| MAPE(%) | 17.63±0.97 | ||||||||||||||||||
+| MRE(%) | 10.78±0.44 | ||||||||||||||||||
+We provide more experimental results on universality (i.e., for 1-day and 2-day scenarios), as shown in Tables C.2 and C.3.
 We observe that while the ExoST exhibits some negative gains in the 1-day scenario, it exhibits improved performance as the prediction length increases, demonstrating the superior generalization of our method.
 The remaining observations show similar analytical results to those in the main text.
-B.2. More Result on Robustness Study
+C.2 More Result on Robustness Study
 | Datasets | AQI-19 | Speed-19 | Intensity-22 | ||||||||||
 | Masking Strategy | MAE | RMSE | MAPE(%) | MRE(%) | MAE | RMSE | MAPE(%) | MRE(%) | MAE | RMSE | MAPE(%) | MRE(%) | |
-| Zero | 20%-Mask | ||||||||||||
-| 40%-Mask | |||||||||||||
-| 60%-Mask | |||||||||||||
-| 80%-Mask | |||||||||||||
-| Random | 20%-Mask | ||||||||||||
-| 40%-Mask | |||||||||||||
-| 60%-Mask | |||||||||||||
-| 80%-Mask | |||||||||||||
-| No Masking | |||||||||||||
+| Zero | 20% | ||||||||||||
+| 40% | 9.23±0.01 | 9.84±0.01 | |||||||||||
+| 60% | |||||||||||||
+| 80% | |||||||||||||
+| Random | 20% | ||||||||||||
+| 40% | 31.40±1.10 | 54.03±1.13 | 11.38±0.40 | ||||||||||
+| 60% | |||||||||||||
+| 80% | 3.78±0.01 | 31.72±0.03 | 9.23±0.01 | 9.84±0.02 | |||||||||
+| No Masking | 9.33±0.13 | 14.24±0.18 | 38.57±0.32 | 36.09±0.49 | 17.630.97 | ||||||||
 | Datasets | AQI-19 | Speed-19 | Intensity-22 | ||||||||||
 | Masking Strategy | MAE | RMSE | MAPE(%) | MRE(%) | MAE | RMSE | MAPE(%) | MRE(%) | MAE | RMSE | MAPE(%) | MRE(%) | |
-| Zero | 20%-Mask | ||||||||||||
-| 40%-Mask | |||||||||||||
-| 60%-Mask | |||||||||||||
-| 80%-Mask | |||||||||||||
-| Random | 20%-Mask | ||||||||||||
-| 40%-Mask | |||||||||||||
-| 60%-Mask | |||||||||||||
-| 80%-Mask | |||||||||||||
-| No Masking | |||||||||||||
-B.3. More Result on Mechanism Study
-B.3.1. Ablation Study.
+| Zero | 20% | 50.06±6.65 | 89.00±13.90 | 26.06±2.03 | 18.12±2.40 | ||||||||
+| 40% | |||||||||||||
+| 60% | |||||||||||||
+| 80% | |||||||||||||
+| Random | 20% | 9.32±0.05 | |||||||||||
+| 40% | |||||||||||||
+| 60% | |||||||||||||
+| 80% | 9.32±0.04 | ||||||||||||
+| No Masking | 10.04±0.19 | 15.12±0.26 | 42.62±0.52 | 38.84±0.75 | 3.80±0.01 | 31.78±0.02 | 9.32±0.03 | 9.89±0.03 | |||||
+C.3 More Result on Mechanism Study
+C.3.1 Ablation Study.
 | P | F | D | MAE | RMSE | MAPE (%) | MRE (%) |
-[…]
+| ✓ | ✓ | - | 42.59±0.92 | |||
+| ✓ | ✓ | ✓ | 10.04±0.19 | 15.12±0.26 | 38.84±0.75 |
 | P | F | U | MAE | RMSE | MAPE (%) | MRE (%) |
 | ✓ | ✓ | - | 43.71±0.70 | |||
-[…]
-From the data perspective, we also evaluated the impact of exogenous variables on ExoST’s performance for 2-day and 3-day forecasting on the AQI-19 , as shown in Tables 6 and Tables 7.
+| ✓ | ✓ | ✓ | 10.41±0.12 | 15.51±0.16 | 40.34±0.47 |
+From the data perspective, we also evaluated the impact of exogenous variables on ExoST’s performance for 2-day and 3-day forecasting on the AQI-19 , as shown in Tables C.7 and Tables C.7.
 We have further observation: ❶ Consistent with the 1-day findings, future covariates outperform past covariates, while data variables alone produce limited accuracy but enhance performance when combined with others, with the best results achieved by integrating all variables, demonstrating ExoST’s ability to synthesize diverse temporal information.
@@
 ❹ Across 1-day to 3-day horizons, the most critical trend is increasing the importance of the past covariates as the forecast horizon extends, complementing the dominant predictive power of future covariates.
-B.3.2. Strategy Study.
+C.3.2 Strategy Study.
 We first formalize the four strategies in detail through formulas:
@@
 The forecast is obtained by averaging the two outputs:
-Simple Weight: Each stream is processed by its own encoder 3.2. The final prediction employs fixed, equal weights:
+Simple Weight: Each stream is processed by its own encoder 4.1. The final prediction employs fixed, equal weights:
 Learnable Weight: Let be a learnable vector, normalized by softmax:
@@
 where the last term is a residual connection.
-Attention: Let denote the intermediate hidden states produced by the two branches before the final read-out layer. For each directional attention we use shared projection matrices .
-Future Past
+Attention: Let denote the intermediate hidden states produced by the two branches before the final read-out layer. For each directional attention we use shared projection matrices . Future Past
 Past Future
 Final Prediction
-As Figures 10 and 11 show, we can observe that: ❶ In the 2-day horizon, ExoST maintains its advantage, while Simple Weight and Learnable Weight exhibit amplified weaknesses: the former’s static weighting fails to handle increased uncertainty in longer forecasts, and the latter, despite its adaptable weights, amplifies noise from future exogenous signals due to lacking day-to-day context modeling.
-❷ Althougn all methods show increasing errors across horizons of 1-day to 3-day, ExoST consistently suppresses error accumulation better than alternatives.
-Appendix C More Disscuison
-C.1. Limitation
+As Figures C.1 show, we can observe that: ❶ In the 2-day horizon, ExoST maintains its advantage, while Simple Weight and Learnable Weight exhibit amplified weaknesses: the former’s static weighting fails to handle increased uncertainty in longer forecasts, and the latter, despite its adaptable weights, amplifies noise from future exogenous signals due to lacking day-to-day context modeling.
+❷ Although all methods show increasing errors across horizons of 1-day to 3-day, ExoST consistently suppresses error accumulation better than alternatives.
+Appendix D More Discussion
+D.1 Current Limitation
 In this paper, we propose a new framework for modeling exogenous variables in spatio-temporal forecasting: ExoST.
@@
 Second, our model currently assumes that spatio-temporal exogenous variables are numerical sensor signals.
-However, real-world scenarios often involve diverse, heterogeneous exogenous information, such as language and visual information.
-Recent research has explored using Large Language Models (LLMs) for time series (jin, 2024; Pan et al., 2024; Sun et al., [n.
-d.]; Jia et al., 2024; Chang et al., 2025; Liu et al., 2025b) and spatio-temporal forecasting (Liu et al., 2025a, 2024; Li et al., 2024; Xu and Liu, 2025), hoping that LLMs can gain numerical prediction capabilities through various alignment techniques.
-However, some studies have raised significant doubts about these approaches (Tan et al., 2024).
+However, real-world scenarios often involve diverse, heterogeneous exogenous information, such as language and visual information (chen2024terra; luo2024towards; chen2025climateiqa).
+Recent research has explored using Large Language Models (LLMs) for time series (jin2024position; pan2024s; suntest; jia2024gpt4mts; chang2025llm4ts; liu2025timecma) and spatio-temporal forecasting (liu2024spatial; li2024urbangpt; liu2025st; xu2025gpt4tfp), hoping that LLMs can gain numerical prediction capabilities through various alignment techniques.
+However, some studies have raised significant doubts about these approaches (tan2024language).
 Different from these views, we believe that language can be more effectively viewed as sparse, discrete exogenous auxiliary information.
-The key challenge lies in how to leverage this information to enhance various spatio-temporal tasks.
-C.2. Future Direction
+The key challenge lies in how to leverage this information to enhance various ST tasks.
+D.2 Future Direction
 Based on these analyses, we identify two key areas for future work:
@@
 This approach is distinct from using language models as the primary prediction backbone.
-For spatio-temporal reasoning tasks, we believe the core essence is to activate the intrinsic capabilities of language models to analyze the underlying causes of spatio-temporal changes and provide logical knowledge-based answers.
+For spatio-temporal reasoning tasks (han2025large; li2025stbench; hettige2025modular; ni2026streasoner), we believe the core essence is to activate the intrinsic capabilities of language models to analyze the underlying causes of ST changes and provide logical knowledge-based answers.